### 05 Jan 2024 v2

# Joint User Association and Resource Allocation for Multi-Cell Networks with Adaptive Semantic Communication

Abstract
Semantic communication is a promising communication paradigm that utilizes Deep Neural Networks (DNNs) to extract the information relevant to downstream tasks, hence significantly reducing the amount of transmitted data. In current practice, the semantic communication transmitter for a specific task is typically pre-trained and shared by all users. However, due to user heterogeneity, it is desirable to use different transmitters according to the available computational and communication resources of users. In this paper, we first show that it is possible to dynamically adjust the computational and communication overhead of DNN-based transmitters, thereby achieving adaptive semantic communication. After that, we investigate the user association and resource allocation problem in a multi-cell network where users are equipped with adaptive semantic communication transmitters. To solve this problem, we decompose it into three subproblems involving the scheduling of each user, the resource allocation of each base station (BS), and the user association between users and BSs. Then we solve each problem progressively based on the solution of the previous subproblem. The final algorithm can obtain near-optimal solutions in polynomial time. Numerical results show that our algorithm outperforms benchmarks under various situations.
Index Terms:
Semantic communication, resource allocation, user association, multi-cell networks
I Introduction
It has been over seventy years since the inception of Shannon’s information theory [1]. During this period, traditional wireless communication technologies, aiming at the accurate transmission of bits, have made significant advancements and approached the Shannon limit. However, with the emergence of modern applications such as virtual reality and smart cities, the amount of data that needs to be transmitted continues to grow at an unprecedented pace. To meet the transmission demands with limited communication resources, a new communication paradigm called semantic communication has been proposed [2, 3, 4]. With the help of artificial intelligence technologies, semantic communication only transmits the information relevant to downstream tasks (which is usually referred to as semantic information), hence leading to a significant reduction in data traffic.
Although semantic communication can reduce the amount of transmitted data, extracting semantic information introduces additional computational overhead. In current practice, users performing the same task typically use the same semantic communication transmitter. However, due to the heterogeneity of users and networks, it is desirable to dynamically adjust the transmitter according to the available computational and communication resources. For example, since the transmitters are implemented by Deep Neural Networks (DNNs), we can decrease the model size of DNNs for users with limited computational capabilities. Similarly, for users with poor channel conditions, we can further reduce the amount of transmitted data by only transmitting a portion of the semantic information. By combining these two methods, we can obtain a range of semantic communication schemes with different computational overhead, communication overhead, and result accuracies. By selecting the optimal semantic communication scheme for each user, we can achieve adaptive semantic communication and significantly improve the system performance.
In this paper, we consider the joint user association and resource allocation problem for multi-cell networks with adaptive semantic communication. However, due to user heterogeneity (e.g. variations in computational capabilities, maximum transmitting power, and energy budget) and network heterogeneity (e.g. variations in user distribution, base station coverage, and channel conditions), there does not exist a straightforward solution for this problem. To solve this problem, we need to address the following challenges.
  • Firstly, both computation and communication processes consume energy and time. Therefore, there exists a trade-off between the computational workload and the communication workload that a user can handle. Since the result accuracy depends on both kinds of workloads, we need to quantify this trade-off in order to obtain the optimal semantic communication scheme. However, how to quantify this trade-off is not explicitly clear.
  • Secondly, in addition to transmitting power and transmission time, the spectrum bandwidth also influences the amount of transmitted data, thereby further affecting the result accuracy. However, due to user heterogeneity, the impact of spectrum bandwidth on result accuracy varies from user to user and hard to quantify. As a result, there is no simple rule to find the optimal spectrum allocation.
  • Thirdly, due to network heterogeneity, associating all users with the nearest base station (BS) may result in severe load imbalance. Therefore, to make appropriate user association decisions, we need to consider both load balancing and channel conditions comprehensively. However, it is not clear how to measure the impact of these two factors.
To address the above challenges, this paper designs an efficient approximate algorithm to obtain near-optimal user association and resource allocation decisions in polynomial time. Our main contributions can be summarized as follows:
  • We propose the concept of adaptive semantic communication and formulate the user association and resource allocation problem in a multi-cell network. We also show the formulated problem is strongly NP-hard.
  • We progressively solve the problem by decomposing it into three subproblems involving the scheduling of each user, the resource allocation of each BS, and the user association between users and BSs. The solution to each subproblem relies on the properties of the optimal solution of the previous subproblem. The final algorithm can obtain near-optimal solutions in polynomial time.
  • We conduct extensive simulations to evaluate the proposed algorithm under various settings. Numerical results demonstrate that our algorithm achieves significant improvement over benchmark algorithms.
The rest of the paper is organized as follows. In Section II, we review related works. In Section III, we describe the system model and formulate our optimization problem. Our algorithm design and theoretical analysis are presented in Section IV. In Section V, we show the numerical results to validate the performance of our algorithm. Section VI concludes the paper and discusses open problems for future work.
II Related Work
User association is an effective approach to balance the workload among multiple BSs, hence improve the overall system performance. In [5], the authors formulate the user association problem as an integer programming prolem and develop an efficient iterative algorithm by relaxing the integral constraints. However, the obtained solution allows users to associate with multiple BSs, which is usually not realistic in practice. The authors in [6] consider the user association problem in the context of massive multiple-input-multiple-output networks. The objective is to maximize the network utility by optimizing the fraction of time that a user associate with each BS. This problem is further extended to the case with cooperative BSs in [7]. Although user association is well studied in the conventional communication paradigm [8, 9, 10], these methods are not applicable in semantic communication because we also need to consider the influence of users’ computational capabilities on their transmission demands.
The concept of semantic communication has been proposed for a long time [11]. Recently, with the advancement of artificial intelligence technology, there has been a growing interest in utilizing DNNs to implement semantic communication. The authors in [12] designed a semantic communication framework for text transmission. Their approach is extended to construct semantic communication systems for other types of sources (e.g. image[13], speech[14], and video[15]), which has shown significant improvements over the conventional communication paradigm. To improve the performance of semantic communication systems, the authors in [16, 17, 18, 19, 20, 21] investigate the optimal resource allocation in various situations. However, they did not take into account the influence of users’ computational capabilities. Besides, most of these works only consider a single BS, which may lead to suboptimal result in ultra-dense networks.
III System Model and Problem Formulation
In this section, we first present the considered system and describe the concept of adaptive semantic communication. After that, we formulate an utility maximization problem and prove it is strongly NP-hard.
III-A Network Model
Figure 1: A simple example of the considered system.
As shown in Fig. 1, we consider a multi-cell network that consists of $M$ BSs and $N$ wireless devices (WDs), where one WD may communicate with multiple BSs. The set of all BSs and WDs are denoted by $\mathcal{M}$ and $\mathcal{N}$, respectively. There are a set $\mathcal{A}$ of $A$ applications running in the system. Let $\mathcal{N}_{a}\subseteq\mathcal{N}$ be the set of WDs belonging to application $a\in\mathcal{A}$. Each WD in $\mathcal{N}_{a}$ has a mount of raw data to be transmitted to BSs for further processing. For example, the WDs may be surveillance cameras that send captured photos to BSs for security risk detection. Due to limited wireless resources, the WDs will first extract the small-sized semantic information from the raw data (e.g. feature maps of the original photos) and only send the extracted semantic information to BSs.
In our considered model, a WD may be within the coverage of multiple BSs, so we need to determine which BS it communicates with. Specifically, the user association is denoted by the binary variable $x^{m}_{n}$, where $x^{m}_{n}=1$ if WD $n$ is associated with BS $m$, and $x^{m}_{n}=0$ otherwise. We assume a WD can communicate with at most one BS so $x^{m}_{n}$ should satisfy
$\sum_{m\in\mathcal{M}}x^{m}_{n}\leq 1,\quad\forall n\in\mathcal{N}.$ (1)
III-B Adaptive Semantic Communication Model
In existing semantic communication works, it is commonly assumed that the same transmitter is shared by all WDs. However, due to user heterogeneity, different WDs may prefer different semantic communication schemes. For example, if a WD has weak computational power but good channel conditions and rich spectrum resources, it may transmit the raw data to the BS without extracting the semantic information. In contrast, if a WD has strong computational power but limited communication resources, it can fully process the raw data and only send the final computation results to the BS. Therefore, by choosing appropriate semantic communication schemes, we can optimize the utilization of communication and computational resources, thereby enhancing system performance.
In practice, there are several ways to generate different semantic communication schemes. For example, the works in [22, 23, 24, 25] proposed various techniques to train DNNs of different model sizes in a single training process. Smaller models require less computation but also have worse accuracy. Since the semantic transmitters are implemented by DNNs, we can use these techniques to obtain a series of transmitters to achieve efficiency-accuracy trade-offs. Another method to derive multiple semantic communication schemes is through semantic compression [18, 19, 20, 21]. The basic idea is, after extracting the semantic information, the transmitter evaluates the importance of each component of the extracted semantic information and only transmit those that contribute most to the result accuracy. The evaluation of semantic information induces extra computation but the amount of data to be transmitted can be significantly reduced.
According to the above discussion, for a given application $a$, its result accuracy is determined by the computation workload $c^{a}$ (defined as the number of CPU cycles required to extract the semantic information) and the communication workload $d^{a}$ (defined as the amount of data to be transmitted). Therefore, we can use the utility function $u^{a}(c,d)$ to represent the obtained utility when the computation and communication workload is $c$ and $d$. It is worth noting that in practice, $c$ and $d$ are often limited to discrete values. For example, in semantic compression, each component of the semantic information cannot be further divided so $d$ must be an integer multiple of the data size of each component. However, the possible values of $c$ and $d$ are usually relatively dense so $u^{a}(c,d)$ can be regarded as a sufficiently good approximation.
III-C Computation and Transmission Model
In current practice, semantic information extraction is usually performed by DNNs, which requires a large amount of computation. For energy efficiency considerations, we assume WDs adopt the Dynamic Voltage and Frequency Scaling (DVFS) technique [26] to dynamically adjust the CPU frequency. For WD $n\in\mathcal{N}_{a}$, let $f_{n}$ be its CPU frequency and $c_{n}$ be its computation workload, then the computation time for semantic information extraction is
$T^{c}_{n}=\frac{c_{n}}{f_{n}}.$ (2)
According to the results in [27], the energy consumption per CPU cycle is proportional to the square of the frequency, which can be expressed as $\gamma_{n}f^{2}_{n}$, where $\gamma_{n}$ is the energy-efficiency factor determined by the chip architecture. Hence, the energy consumption due to computation is
$E^{c}_{n}=c_{n}\gamma_{n}f^{2}_{n}.$ (3)
The spectrum bandwidth of each BS is equally divided into resource blocks (RBs) with bandwidth $W$ and the number of available RBs at BS $m$ is $K_{m}$. Let $z^{m}_{n}\in\{0,1,\dots,K_{m}\}$ be the number of BS $m$’s RBs allocated to WD $n$. As in [28, 29, 30, 31], we assume the inter-BS interference is randomized and/or mitigated by various coordination techniques so the total interference seen by each BS can be regarded as static. According to the Shannon capacity formula, the transmission rate from WD $n$ to BS $m$ can be expressed as:
$r^{m}_{n}=z^{m}_{n}W\log_{2}\left(1+\frac{P_{n}h^{m}_{n}}{\sigma^{2}+I_{m}}\right),$ (4)
where $P_{n}$ is the transmission power of WD $n$, $h^{m}_{n}$ is the channel gain between WD $n$ and BS $m$, $\sigma^{2}$ is the noise power at BS $m$, and $I_{m}$ is the average interference seen by BS $m$. Clearly, for each BS $m$, the number of totally allocated RBs cannot exceed $K_{m}$, so we have
$\sum_{n\in\mathcal{N}}z^{m}_{n}\leq K_{m},\quad\forall m\in\mathcal{M}.$ (5)
For WD $n\in\mathcal{N}_{a}$, let $d_{n}$ be the amount of data to be transmitted, then the transmission time is
$T^{t}_{n}=\frac{d_{n}}{\sum_{m\in\mathcal{M}}x^{m}_{n}r^{m}_{n}}$ (6)
and the energy consumption due to transmission is
$E^{t}_{n}=P_{n}T^{t}_{n}.$ (7)
III-D Problem Formulation
In this paper, we aim to maximize the total utility by jointly optimizing the user association and resource allocation under energy consumption, wireless bandwidth, and completion time constraints. Based on the models described above, the corresponding optimization problem can be formulated as follows:
$\max_{\begin{subarray}{c}\bm{x},\bm{z},\bm{f},\bm{P}\\ \bm{c},\bm{d}\end{subarray}}\quad$ $\sum_{a\in\mathcal{A}}\sum_{n\in\mathcal{N}_{a}}u^{a}(c_{n},d_{n})$ (10)
[…]
$T^{c}_{n}+T^{t}_{n}\leq T^{max},\quad\forall n\in\mathcal{N}$ (10b)
$E^{c}_{n}+E^{t}_{n}\leq E^{max}_{n},\quad\forall n\in\mathcal{N}$ (10c)
$0\leq f_{n}\leq f^{max}_{n},\quad\forall n\in\mathcal{N}$ (10d)
$0\leq P_{n}\leq P^{max}_{n},\quad\forall n\in\mathcal{N}$ (10e)
$x^{m}_{n}\in\{0,1\},\quad\forall m\in\mathcal{M},\forall n\in\mathcal{N}$ (10f)
$z^{m}_{n}\in\{0,1,\dots,K_{m}\},\quad\forall m\in\mathcal{M},\forall n\in\mathcal{N}$ (10g)
where $\bm{x},\bm{z},\bm{f},\bm{P},\bm{c},\bm{d}$ are vetors of corresponding decision variables. Constraint (10a) contains the user association and RB allocation constraints described above. Constraint (10b) ensures the semantic information transmission is completed before the maximum delay $T^{max}$. Constraint (10c) requires the total energy consumption of WD $n$ does not exceed a given threshold $E^{max}_{n}$. Constraints (10d), (10e), (10f), and (10g) define the ranges of feasible values for the decisions variables, where $f^{max}_{n}$ and $P^{max}_{n}$ are the maximum CPU frequency and transmission power of WD $n$, respectively.
The formulated problem is Mixed-Integer Programming (MIP) and can be hard to solve. Specifically, we will show that our problem is strongly NP-hard, so it does not even have a pseudo-polynomial time algorithm. Before stating our result, we give the definition of binary utility functions: we say an utility function $u^{a}(c,d)$ is binary if $u^{a}(c,d)=u^{a}_{0}$ for all $(c,d)\geq(c^{a}_{0},d^{a}_{0})$ and $u^{a}(c,d)=0$ otherwise, where $u^{a}_{0},c^{a}_{0},d^{a}_{0}$ are positive constants.
Theorem 1.
Problem (10) is strongly NP-hard even if (i) there is only one WD for each running application and its utility function is binary, (ii) the energy budget $E^{max}_{n}$ is sufficiently large, and (iii) the channel gain $h^{m}_{n}$ and interference $I_{m}$ for different BS $m$ is equal.
Proof.
Please see Appendix A. ∎
Thereom 1 indicates that problem (10) is hard to solve even for a significantly simplified case. Therefore, in general, it is computationally intractable to obtain the exact solution of problem (10). In the next section, we will design an approximate algorithm that produces near-optimal solutions for problem (10) in polynomial time.
IV Algorithm Design and Performance Analysis
In this section, we will progressively solve the formulated problem through three stages. Firstly, we find the optimal scheduling decisions $\bm{f},\bm{P},\bm{c},\bm{d}$ when the user association and RB allocation are given. Based on this result, we can solve the optimal RB allocation under any user association. Finally, we can obtain the optimal user association by employing a relaxation-then-refinement approach.
IV-A Scheduling Subproblem for Each WD
If the user association $\bm{x}$ and RB allocation $\bm{z}$ are given, problem (10) can be decomposed into $N$ subproblems concerning the scheduling of each WD. Specifically, for any WD $n\in\mathcal{N}_{a}$ associated with BS $m$, its scheduling subproblem can be formulated as:
$\max_{{f_{n}},{P_{n}},c_{n},d_{n}}\quad$ $u^{a}(c_{n},d_{n})$ (11)
$s.t.\quad$ $T^{c}_{n}+T^{t}_{n}\leq T^{max},$ (11a)
$E^{c}_{n}+E^{t}_{n}\leq E^{max}_{n},$ (11b)
$0\leq f_{n}\leq f^{max}_{n},0\leq P_{n}\leq P^{max}_{n}$ (11c)
We say a computation and communication workload pair $(c_{n},d_{n})$ is achievable if there exists $f_{n},P_{n}$ such that constraints (11a)-(11c) are satisfied. The set of all achievable workload pairs is referred to as the achievable region. Since $u^{a}(c_{n},d_{n})$ is non-decreasing, the optimal workload pair $(c^{*}_{n},d^{*}_{n})$ must lie on the boundary of the achievable region, i.e. there is no achievable $(c_{n},d_{n})$ such that $c_{n}=c^{*}_{n}$ and $d_{n}>d^{*}_{n}$. To obtain the achievable region, we only need to calculate the maximum achievable $d_{n}$ for any given $c_{n}$, denoted by $d^{max}_{n}(c_{n})$. The calculation of $d^{max}_{n}(c_{n})$ can be formulated as another optimization problem
$\max_{f_{n},P_{n}}\quad$ $d_{n}$ (12)
[…]
To solve problem (12), we first show that its optimal solution satisfies the following condition.
Theorem 2.
The optimal solution $(f^{*}_{n},P^{*}_{n})$ of problem (12) must satisfy
$T^{c}_{n}+T^{t}_{n}$ $=T^{max}$ (13)
$E^{c}_{n}+E^{t}_{n}$ $=E^{max}_{n}.$ (14)
Proof.
Please see Appendix B. ∎
According to Theorem 2, we can replace constraint (11a) and (11b) with (13) and (14). Substituting (2) and (6) into (13) and recalling that WD $n$ is associated with BS $m$ (i.e. $x^{m}_{n}=1$) result in
$P_{n}=\left(2^{\frac{d_{n}f_{n}}{(f_{n}T^{max}-c_{n})z^{m}_{n}W}}-1\right)\frac{\sigma^{2}+I_{m}}{h^{m}_{n}}.$ (15)
Combining (15) with (11c) we get the feasible region of $f_{n}$:
$\frac{z^{m}_{n}Wc_{n}\log_{2}\left(1+\frac{P_{n}^{max}h^{m}_{n}}{\sigma^{2}+I_{m}}\right)}{z^{m}_{n}WT^{max}\log_{2}\left(1+\frac{P_{n}^{max}h^{m}_{n}}{\sigma^{2}+I_{m}}\right)-d_{n}}\leq f_{n}\leq f_{n}^{max}.$ (16)
Substituting (3), (7) and (15) into (14) yields
$d_{n}=$ $z^{m}_{n}W\left(T^{max}-\frac{c_{n}}{f_{n}}\right)\times$
$\quad\log_{2}\left(1+\frac{h^{m}_{n}(E^{max}-c_{n}\gamma_{n}f^{2}_{n})}{(T^{max}-\frac{c_{n}}{f_{n}})(\sigma^{2}+I_{m})}\right).$ (17)
Hence, problem (12) is equivalent to finding the optimal $f_{n}$ that maximizes (17) subject to (16), which can be solved by letting $\frac{\partial d_{n}}{\partial f_{n}}=0$ and comparing the values of $d_{n}$ achieved by every critical point in the feasible region. Clearly, the optimal $f_{n}$ is a function of $c_{n}$, denoted by $f^{*}_{n}(c_{n})$. According to (17), we can regard $d_{n}$ as a function of $c_{n}$ and $f_{n}$. Then we have $d^{max}_{n}(c_{n})=\max_{f_{n}}d_{n}(c_{n},f_{n})=d_{n}(c_{n},f^{*}_{n}(c_{n}))$. Unfortunately, the expression of $\frac{\partial d_{n}}{\partial f_{n}}=0$ is fairly complex so we are unable to derive an analytical expression for $f^{*}_{n}(c_{n})$. However, we can still calculate the value of $f^{*}_{n}(c_{n})$ through numerical methods.
Once we know how to calculate $d^{max}_{n}(c_{n})$, problem (11) is equivalent to $\max_{c_{n}}u^{a}(c_{n},d^{max}_{n}(c_{n}))$. Similarly, the optimal $c_{n}$ can be obtained by letting
$\frac{\mathop{}\!\mathrm{d}u^{a}(c_{n},d^{max}_{n}(c_{n}))}{\mathop{}\!\mathrm{d}c_{n}}$
$\quad=\frac{\partial u^{a}}{\partial c_{n}}+\frac{\partial u^{a}}{\partial d^{max}_{n}}\times\frac{\mathop{}\!\mathrm{d}d^{max}_{n}}{\mathop{}\!\mathrm{d}c_{n}}$
$\quad=\frac{\partial u^{a}}{\partial c_{n}}+\frac{\partial u^{a}}{\partial d^{max}_{n}}\times\left(\frac{\partial d_{n}}{\partial c_{n}}+\frac{\partial d_{n}}{\partial f^{*}_{n}}\times\frac{\mathop{}\!\mathrm{d}f^{*}_{n}}{\mathop{}\!\mathrm{d}c_{n}}\right)$
$\quad=\frac{\partial u^{a}}{\partial c_{n}}+\frac{\partial u^{a}}{\partial d^{max}_{n}}\times\frac{\partial d_{n}}{\partial c_{n}}=0$
where the second equality holds because $d^{max}_{n}(c_{n})=d_{n}(c_{n},f^{*}_{n}(c_{n}))$ and the third equality holds because $\frac{\partial d_{n}}{\partial f^{*}_{n}}=0$. Similarly, the optimal $c_{n}$ can only be solved by numerical methods. After obtaining the optimal $c_{n}$, the optimal values of the rest decision variables can be calculated easily.
IV-B RB Allocation Subproblem for Each BS
In the previous subsection, we have obtained the optimal $\bm{f},\bm{P},\bm{c},\bm{d}$ for any given $\bm{x}$ and $\bm{z}$. In this subsection, we still assume $\bm{x}$ is given and show how to find the optimal $\bm{z}$. Let $u^{*}_{n}(z^{m}_{n})$ be the optimal objective value of problem (11). The meaning of $u^{*}_{n}(z^{m}_{n})$ is the maximum achievable utility of WD $n$ when $z^{m}_{n}$ RBs are allocated to it. Since $\bm{x}$ is given, we can use $\mathcal{N}_{m}$ to denote the set of WDs associated with BS $m$. Then the RB allocation subproblem for BS $m$ becomes
$\max_{z^{m}_{n}}\quad$ $\sum_{n\in\mathcal{N}_{m}}u^{*}_{n}(z^{m}_{n})$ (18)
$s.t.\quad$ $\sum_{n\in\mathcal{N}_{m}}z^{m}_{n}\leq K_{m},$ (18a)
$z^{m}_{n}\in\{0,1,\dots,K_{m}\},\quad\forall n\in\mathcal{N}_{m}.$ (18b)
In this subsection, we will design two different algorithms for problem (18) based on the properties of the utility function $u^{a}(c,d)$. In practice, the most commonly used utility function is the result accuracy, e.g. the classification accuracy in image recognition tasks. In this case, the utility function usually exhibits concavity and other properties that can be exploited to accelerate the solving of problem (18). However, in some scenarios, we may encounter more general utility functions. Therefore, we also design an algorithm that does not require the concavity of the utility function.
IV-B1 RB Allocation with Concave Utility
If the utility function is the result accuracy or a concave function of the result accuracy, then it should satisfy the following three properties. Firstly, according to the results in [22, 23, 24, 25, 18, 19, 20, 21], the marginal return of $c$ and $d$ is diminishing. Hence, we can assume:
Assumption 1.
$u^{a}(c,d)$ is non-descreasing and concave.
Secondly, the marginal return of computation workload should be non-decreasing with respect to the amount of transmitted data. This is because the improvement in result accuracy due to increased computation needs to be propagated to the destination through data transmission. If the amount of transmitted data is limited, then the enhancement obtained at the destination will be discounted. An extreme example is the marginal return of $c$ is 00 when $d=0$, as $u^{a}(c,0)=0$ for all $c$. Therefore, we have:
Assumption 2.
$\frac{\partial u^{a}}{\partial c}$ is non-decreasing with respect to $d$.
Thirdly, according to the results in [19], the marginal return of $d$ drops rapidly, which means the partial derivative of $u^{a}(c,d)$ with respect to $d$ also decreases rapidly. Hence, we can assume $\frac{\partial u^{a}}{\partial d}$ satisfies the following property:
Assumption 3.
$\frac{\partial u^{a}}{\partial d}|_{(c,\alpha d)}\leq\frac{1}{\alpha}\frac{\partial u^{a}}{\partial d}|_{(c,d)}$ for any $\alpha\geq 1$.
Assumption 3 requires $\frac{\partial u^{a}}{\partial d}$ decreases at least at a linear rate of $d$. A common function that satisfies this assumption is the logarithmic function. In practice, this assumption generally does not hold strictly, but it can be regarded as approximately valid within the range of values that we are interested in.
With the above assumptions, we can proceed to demonstrate the concavity of $u^{*}_{n}(z^{m}_{n})$. However, $z^{m}_{n}$ is a discrete variable and the concavity for discrete functions is not well-defined. Therefore, we relax $z^{m}_{n}$ to a real-value variable $\hat{z}^{m}_{n}$ and prove that $u^{*}_{n}(\hat{z}^{m}_{n})$ is concave.
Theorem 3.
The function ${u}^{*}_{n}(\hat{z}^{m}_{n})$ is concave.
Proof.
Please see Appendix C. ∎
Due to the concavity of $u^{*}_{n}(\hat{z}^{m}_{n})$, to solve (18), we only need to greedily allocate RBs to the WD that contributes most to the increase of utility. The detailed steps are as follows. At the beginning, we initialize $z^{m}_{n}=0$ for all $n\in\mathcal{N}_{m}$. After that, for each WD $n\in\mathcal{N}_{m}$, we calculate the increase of utility if an extra RB is allocated to it, i.e. $\Delta u_{n}=u^{*}_{n}(z^{m}_{n}+1)-u^{*}_{n}(z^{m}_{n})$. If WD $n^{\prime}$ has the greatest increase of utility, i.e. $n^{\prime}=\operatorname*{arg\,max}_{n\in\mathcal{N}_{m}}\Delta u_{n}$, we allocate one extra RB to it and update $z^{m}_{n^{\prime}}=z^{m}_{n^{\prime}}+1$. This process continues until all of the $K_{m}$ RBs are allocated.
Our algorithm is summarized in Algorithm 1. Notice that we only modify the values of $z^{m}_{n}$ for one WD at each iteration. Therefore, for the loop in line $3$-$5$, we only need to calculate $\Delta u_{n}$ for all $n\in\mathcal{N}_{m}$ for the first time we entered the loop. After that, we only need to update $\Delta u_{n}$ for the WD whose $z^{m}_{n}$ is modified in the previous iteration. Apparently, the time complexity of the whole algorithm is $O(|\mathcal{N}_{m}|K_{m})$, where $|\mathcal{N}_{m}|$ is the number of elements in $\mathcal{N}_{m}$.
Algorithm 1 Algorithm for the RB Allocation Subproblem with Concave Utility
1:  Initialization: $z^{m}_{n}=0,\forall n\in\mathcal{N}_{m}$;
2:  for $i=\{1,2,\dots,K_{m}\}$ do
3:     for $n\in\mathcal{N}_{m}$ do
4:        Calculate $\Delta u_{n}=u^{*}_{n}(z^{m}_{n}+1)-u^{*}_{n}(z^{m}_{n})$;
5:     end for
6:     Find $n^{\prime}=\operatorname*{arg\,max}_{n\in\mathcal{N}_{m}}\Delta u_{n}$;
7:     Update $z^{m}_{n^{\prime}}=z^{m}_{n^{\prime}}+1$;
8:  end for
9:  return  $z^{m}_{n}$.
IV-B2 RB Allocation with General Utility
For general utility functions, we cannot leverage the concavity of $u^{*}_{n}(\hat{z}^{m}_{n})$ to facilitate algorithm design. However, we can still solve the RB allocation subproblem (18) through dynamic programming. Specifically, let $u_{m}(k,j)$ be the maximum total utility of BS $m$ if there are only $k$ available RBs and $j$ WDs in BS $m$, i.e. $u_{m}(k,j)$ is the optimal objective value of the following optimization problem:
$\max_{z^{m}_{n}}\quad$ $\sum_{n=1}^{j}u^{*}_{n}(z^{m}_{n})$ (19)
$s.t.\quad$ $\sum_{n=1}^{j}z^{m}_{n}\leq k$
$z^{m}_{n}\in\{0,1,\dots,k\},\quad\forall n\in\{1,\dots,j\}.$
Apparently, the optimal objective value of problem (18) equals $u_{m}(K_{m},|\mathcal{N}_{m}|)$. For convenience, we use $z_{m}(k,j)$ to denote the optimal RB allocation for the last WD (i.e. WD $j$) in (19). Clearly, when there is only one WD in the system, we should allocate all RBs to it. Therefore, we have
$u_{m}(k,1)=u^{*}_{1}(k)\mbox{ and }z_{m}(k,1)=k.$
Otherwise, for any $j\in\{2,3,\dots,N\}$, the value of $u_{m}(k,j)$ can be calculated through
$u_{m}(k,j)=\max_{z^{m}_{j}}\{u_{m}(k-z^{m}_{j},j-1)+u^{*}_{j}(z^{m}_{j})\}.$ (20)
Correspondingly, $z_{m}(k,j)$ is the optimal $z^{m}_{j}$ that maximizes (20). Based on (20), we can finally obtain $u_{m}(K_{m},|\mathcal{N}_{m}|)$. Then the optimal $z^{m}_{n}$ can be obtained by backtracking the optimal allocation decisions recorded in $z_{m}(k,j)$. Specifically, we can let $z^{m}_{|\mathcal{N}_{m}|}=z_{m}(K_{m},|\mathcal{N}_{m}|)$ and $z^{m}_{j}=z_{m}(K_{m}-\sum_{n=j+1}^{|\mathcal{N}_{m}|}z^{m}_{n},j)$ for $j\in\{1,\dots,|\mathcal{N}_{m}|-1\}$.
The detailed steps are summarized in Algorithm 2. When calculating $u_{m}(k,j)$ using (20), we need to iterate over all possible values of $z^{m}_{j}$, which amount to a total of $k$ values. Then one can easily verify that the time complexity of Algorithm 2 is $O\left(|\mathcal{N}_{m}|K^{2}_{m}\right)$.
Algorithm 2 Algorithm for the RB Allocation Subproblem with General Utility
1:  Initialization: $u_{m}(k,1)=u^{*}_{1}(k),\ z_{m}(k,1)=k,\ \forall k\in\{1,\dots,K_{m}\}$;
2:  for $j\in\{2,3,\dots,|\mathcal{N}_{m}|\}$ do
3:     for $k\in\{0,1,\dots,K_{m}\}$ do
4:        Calculate $u_{m}(k,j)$ according to (20) and record the corresponding $z_{m}(k,j)$.
5:     end for
6:  end for
7:  Let $k=K_{m}$;
8:  for $j=\{|\mathcal{N}_{m}|,|\mathcal{N}_{m}|-1,\dots,1\}$ do
9:     Let $z^{m}_{j}=z_{m}(k,j)$;
10:     Let $k=k-z^{m}_{j}$;
11:  end for
12:  return  $z^{m}_{n}$.
IV-C User Association Subproblem
In this subsection, we aim to find the optimal user association decision $\bm{x}^{*}$. We first assume that each WD is associated with all BSs, i.e. $\mathcal{N}_{m}=\mathcal{N}$ for all $m\in\mathcal{M}$. By using algorithms developed in the previous subsections, we can calculate the optimal RB allocation and scheduling decisions. The resulting total utility is an upper bound for the objective value of the original problem (10). After that, we progressively transition to a user association decision that satisfies constraint (1) through $N$ steps. Specifically, in each step, we select a WD that does not satisfy constraint (1) and associate it with exactly one BS. Our objective is to minimize the decrease in utility at each step. The remaining problem is: (i) how to select a WD at each step and (ii) how to choose the target BS for the selected WD.
We first address the second problem. Suppose we have selected a WD $n$ that is associated with all BSs. For an arbitrary BS $m$, let $(z^{m}_{n^{\prime}})_{n^{\prime}\in\mathcal{N}_{m}}$ be the optimal RB allocation. If WD $n$ is no longer associated with BS $m$, the RBs it occupies will be released and allocated to other WDs. The new optimal RB allocatation is denoted by $(\bar{z}^{m}_{n^{\prime}})_{n^{\prime}\in\mathcal{N}_{m}\setminus\{n\}}$. Then the decrease in utility for BS $m$ is
$\Delta_{m}=u^{*}_{n}(z^{m}_{n})-\sum_{n^{\prime}\in\mathcal{N}_{m}\setminus\{n\}}\left(u^{*}_{n^{\prime}}(\bar{z}^{m}_{n^{\prime}})-u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})\right).$
Recall that our objective is to minimize the decrease in total utility. Therefore, we only need to associate WD $n$ with the BS $m$ that has the greatest $\Delta_{m}$. When BS $m$ is associated with multiple WDs, the difference between $z^{m}_{n^{\prime}}$ and $\bar{z}^{m}_{n^{\prime}}$ is relatively small. Therefore, we can use the derivative of $u^{*}_{n^{\prime}}$ at $z^{m}_{n^{\prime}}$ to estimate the difference between $u^{*}_{n^{\prime}}(\bar{z}^{m}_{n^{\prime}})$ and $u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})$, i.e. $u^{*}_{n^{\prime}}(\bar{z}^{m}_{n^{\prime}})-u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})\approx\frac{\mathop{}\!\mathrm{d}u^{*}_{n^{\prime}}}{\mathop{}\!\mathrm{d}z^{m}_{n^{\prime}}}\times(\bar{z}^{m}_{n^{\prime}}-z^{m}_{n^{\prime}})$. If $z^{m}_{n^{\prime}}$ is continuous, then according to the KKT conditions, the derivative $\frac{\mathop{}\!\mathrm{d}u^{*}_{n^{\prime}}}{\mathop{}\!\mathrm{d}z^{m}_{n^{\prime}}}$ is equal for all $n^{\prime}$. In our problem, although $z^{m}_{n^{\prime}}$ takes discrete values, the values of derivatives should be close so we can regard them as approximately equal. Therefore, we only need to calculate the derivative for one $n^{\prime}$. Since it is not very convenient to calculate the derivatives for discrete variables, we use the marginal gain $\delta_{m}=u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}}+1)-u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})$ to replace the role of derivatives. Then we have
$\Delta_{m}$ $\approx u^{*}_{n}(z^{m}_{n})-\sum_{n^{\prime}\in\mathcal{N}_{m}\setminus\{n\}}\delta_{m}(\bar{z}^{m}_{n^{\prime}}-z^{m}_{n^{\prime}})$
$=u^{*}_{n}(z^{m}_{n})-\delta_{m}z^{m}_{n}$ (21)
where the equality holds because $\sum_{n^{\prime}\in\mathcal{N}_{m}\setminus\{n\}}\bar{z}^{m}_{n^{\prime}}=K_{m}$ and $\sum_{n^{\prime}\in\mathcal{N}_{m}\setminus\{n\}}{z}^{m}_{n^{\prime}}=K_{m}-z^{m}_{n}$.
The next problem is how to select the appropriate WD at each step. As described above, for any given WD $n$, we choose its target BS based on $\Delta_{m}$. However, the value of $\Delta_{m}$ depends on the set of WDs associated with BS $m$, which is $\mathcal{N}_{m}$. As our algorithm progresses, $\mathcal{N}_{m}$ also changes. Therefore, the BS previously chosen for WD $n$ may no longer be the optimal one. We refer to this phenomenon as “regret”. Intuitively, if we process the WDs in an appropriate order, the resulting regret will be relatively less.
As mentioned above, regret occurs when the values of $\Delta_{m}$ change so that the previously maximum $\Delta_{m}$ no longer be the maximum. Notice that the values of $\Delta_{m}$ change in a gradual manner as $\mathcal{N}_{m}$ varies (i.e. $\Delta_{m}$ will not change abruptly), hence we have the following two observations. Firstly, for any given WD $n$, if the $\Delta_{m}$ for one BS $m$ is significantly higher than the rest BSs, then it is highly probable that BS $m$ is still the optimal choice for WD $n$ even if $\mathcal{N}_{m}$ changes. Secondly, if WD $n$ is processed relatively late, then the subsequent changes in $\mathcal{N}_{m}$ are limited. This implies that the changes in the values of $\Delta_{m}$ are also limited, hence it is unlikely to result in regret. Based on the above observations, to minimize regret, we should defer the processing of WDs whose values of $\Delta_{m}$ are close. To measure the degree of difference among $\Delta_{m}$, we define the kurtosis of WD $n$ as the ratio of the maximum $\Delta_{m}$ to the sum of all $\Delta_{m}$, i.e. $\kappa_{n}=\frac{\max_{m}\Delta_{m}}{\sum_{m}\Delta_{m}}$. Clearly, if the values of $\Delta_{m}$ are close, then the value of $\kappa_{n}$ is relatively small. To defer the processing of WDs with small $\kappa_{n}$, we only need to pick the WD with the largest $\kappa_{n}$ in each step.
The complete algorithm is summarized in Algorithm 3, where $\mathcal{N}_{r}$ denotes the set of WDs to be processed. Let $K=\max_{m}K_{m}$ be the maximum number of RBs in all BSs. Based on the result in the previous subsection, the time complexity of line $3$ is $O(MNK)$ or $O(MNK^{2})$, depending on whether the utility function is concave. Since the time complexity of the loop in line $4$-$9$ is only $O(NM)$, the time complexity of the whole algorithm is $O(MN^{2}K)$ for concave utility functions and $O(MN^{2}K^{2})$ for general utility functions.
Algorithm 3 Algorithm for the User Association Subproblem
1:  Initialization: $\mathcal{N}_{r}=\mathcal{N},x^{m}_{n}=1,\forall n\in\mathcal{N},m\in\mathcal{M}$;
2:  for $i\in\{1,\dots,N\}$ do
3:     Solve the RB allocation subproblem for each $m\in\mathcal{M}$;
4:     for $n\in\mathcal{N}_{r}$ do
5:        for $m\in\mathcal{M}$ do
6:           Estimate $\Delta_{m}$ according to (21);
7:        end for
8:        Calculate $\kappa_{n}=\frac{\max_{m}\Delta_{m}}{\sum_{m}\Delta_{m}}$;
9:     end for
10:     Find $n^{\prime}=\operatorname*{arg\,max}_{n}\kappa_{n}$ and $m^{\prime}=\operatorname*{arg\,max}_{m}\Delta_{m}$;
11:     Set $x^{m^{\prime}}_{n^{\prime}}=1$ and $x^{m}_{n^{\prime}}=0$ for all $m\neq m^{\prime}$;
12:     Update $\mathcal{N}_{r}=\mathcal{N}_{r}\setminus\{n^{\prime}\}$;
13:  end for
14:  return  $x^{m}_{n}$.
V Numerical Results
TABLE I: Simulation Parameters
Parameter Value
Number of RBs in each BS $K_{m}$ $\mathcal{U}\{25,50,75,100\}$
Bandwidth of each RB $W$ $0.2$ MHz
Delay requirement $T^{max}$ $10$ ms
Energy budget of WDs $E^{max}_{n}$ $2$ mJ
Maximum CPU frequency $f^{max}_{n}$ $\mathcal{U}[1,3]$ GHz
Maximum transmission power $P^{max}_{n}$ $0.2$ W
Energy efficiency of WDs’ processors $\gamma_{n}$ $\mathcal{U}[10^{-28},10^{-27}]$
Noise power at BSs $\sigma^{2}$ $10^{-13}$
Average interference at BSs $I_{m}$ $\mathcal{U}[10^{-13},10^{-12}]$
Figure 2: Utility of each BS under the concave utility function.
Figure 3: Utility of each BS under the general utility function.
Figure 2: Utility of each BS under the concave utility function.
Figure 3: Utility of each BS under the general utility function.
Figure 4: Total utility of different algorithms.
Figure 5: Total utility under different numbers of BSs.
Figure 6: Total utility under different numbers of WDs.
Figure 5: Total utility under different numbers of BSs.
Figure 6: Total utility under different numbers of WDs.
Figure 7: Total utility under different delay requirements.
Figure 8: Total utility under different energy budgets.
In this section, extensive simulations are conducted to evaluate the performance of our algorithm. We consider a wireless network consisting of $M=5$ BSs and $N=30$ WDs that are randomly located in a $\times 500$m square area. For the channel model, we assume the pathloss is $128.1+37.6\log_{10}Dist$ dB and the shadowing factor is $6$ dB, where $Dist$ is the distance (in km) between WDs and BSs.
Our utility function is based on the empirical results about the relationship between result accuracy and computational or communication workload. To model the impact of computational workload $c$ on result accuracy, we conduct curve-fitting on the results given in [23], which shows the accuracy of various models under different computational overhead. We find that the accuracy $A_{c}$ can be well approximated by the following logarithmic function $A_{c}=\eta^{a}_{1}log(\frac{c}{C^{a}})+\eta^{a}_{2}$, where $\eta^{a}_{1},\eta^{a}_{2}$ are parameters corresponding to application $a$ and $C^{a}$ is the computational workload of the maximum DNN model. In our simulation, we set $\eta^{a}_{1}\in\mathcal{U}[0.05,0.08],\eta^{a}_{2}\in\mathcal{U}[0.9,0.95]$, and $C^{a}\in\mathcal{U}[5M,10M]$ cycles, where $\mathcal{U}[a,b]$ represents the uniform distribution on interval $[a,b]$.
According to [19], the relationship between result accuracy $A_{d}$ and communication workload $d$ can be modeled by the following function: $A_{d}=\beta^{a}_{1}(1-\frac{d}{D^{a}})^{\beta^{a}_{2}}+\beta^{a}_{3}$, where $\beta^{a}_{1},\beta^{a}_{2},\beta^{a}_{3}$ are parameters and $D^{a}$ is the maximum data size of the DNN outputs. In our simulation, we set $\beta^{a}_{1}\in\mathcal{U}[-0.75,-0.6],\beta^{a}_{2}\in\mathcal{U}[10,20],\beta^{a}_{3}\in\mathcal{U}[0.9,0.95]$ and $D^{a}\in\mathcal{U}[0.15M,0.25M]$ bits. Intuitively, $\beta^{a}_{3}$ is the result accuracy if all DNN outputs are transmitted to the receiver. Hence, the accuracy discount due to limited data transmission is $A_{d}/\beta^{a}_{3}$. Therefore, we assume the final result accuracy at the receiver side is $A=A_{c}\times A_{d}/\beta^{a}_{3}$. To evaluate our algorithm comprehensively, we consider a concave utility function $u^{a}(c,d)=A$ and a general non-concave function $u^{a}(c,d)=\frac{1}{1-A}$. The values of rest parameters are listed in Table I.
To demonstrate the effectiveness of the proposed algorithm (labeled as “Prop”), we compare it with the following benchmarks. Notice that our algorithm consists of three stages. Unless otherwise specified, each benchmark only differs from our algorithm in one stage.
  • Traditional Communication (TC): the raw data is transmitted to the BS without processing.
  • Fixed Semantic Communication (FSC): each WD adopts a fixed semantic communication scheme.
  • Average RB Allocation (ARB): the RBs are allocated to each associated WD evenly.
  • Nearest User Association (NUA): the WDs are associated with the nearest BSs.
  • FSC+ARB+NUA (FAN): we use the above three algorithms to produce decisions in each stage respectively.
For TC, we assume the size of raw data satisfies $D_{raw}\in\mathcal{U}[0.4M,0.8M]$ bits. If WD $n$ successfully transmits the raw data under the delay and energy constraints, then the BS can use the maximum DNN model to process the raw data, which leads to result accuracy $\eta^{a}_{2}$. For FSC, the utilized fixed semantic communication scheme is $c=D^{a}/2$ and $d=D^{a}/2$. The data presented in this section are the average of ten repeated experiments with different random seeds.
In Fig. 4 and Fig. 4, we show the utility of each BS under the concave and general utility functions, respectively. In both cases, our algorithm achieves the largest utility at all BSs, except for BS $5$, where NUA outperforms our algorithm. This is because there are more WDs around BS $5$ and they are all associated with BS $5$ in NUA, hence leading to a high utility. In our algorithm, however, we will associate part of these WDs with other BSs to achieve better load balance. As a result, the total utility of all BSs (which is shown in Fig. 4) produced by our algorithm is higher than that of NUA. Another algorithm that has comparable performance is ARB. This is not surprising because the marginal gain of result accuracy drops very quickly with the increase of $d$. Consequently, the marginal gain of utility also drops rapidly with the growth of allocated RBs. This guarantees the average allocation of RBs can achieve close-to-optimal results.
Compared to the previous algorithms, the performance of TC and FSC degrades significantly. The problem in TC and FSC is similar: they both adopt fixed transmission schemes so they are unable to efficiently utilize resources when dealing with heterogeneous users and networks. An interesting observation is that the performance of FAN is even significantly worse than TC. This indicates that adopting semantic communication alone does not necessarily enhance system performance if the resource allocation is not optimized. According to Fig. 4, most algorithms exhibit similar performance under the two different utility functions, except for TC. This is because, under the general utility function, the utility is the reciprocal of one minus accuracy, hence the utility increases very quickly when the accuracy approaches one. This is beneficial for TC as it achieves particularly high accuracy when a WD successfully transmits the raw data to the BS.
To understand the impact of network scale on algorithm performance, Fig. 7 and Fig. 7 demonstrate the total utility of all algorithms under different numbers of BSs and WDs. Due to space limitations, we only present the data under the concave utility function, but the result is similar under the general utility function. From Fig. 7, one can observe that the utility of all algorithms increases with the number of BSs. However, when the number of BSs exceeds 6, the utility gains for Prop, ARB, and NUA are very limited. In contrast, the algorithms that adopt fixed transmission schemes maintain a relatively high growth rate. However, in Fig. 7, the growth trend is exactly the opposite. Specifically, the total utility of Prop, ARB, and NUA grows continuously, indicating their ability to serve more WDs with existing communication resources. The only algorithm showing a decreased utility is FAN. This is because FAN adopts the average allocation of RBs, so the number of RBs allocated to each WD becomes less when there are more WDs. Since FAN also employs a fixed transmission scheme, some WDs are unable to complete the transmission under increasingly less RBs, hence resulting in zero utility.
In addition to network scale, we also investigated the impact of delay and energy constraints on the algorithm results, as shown in Fig. 7 and Fig. 8. The algorithms most sensitive to delay are FSC and FAN, because an increase in delay allows WDs to process data in a more energy-efficient manner, thereby improving the probability of successful transmission. The impact of energy budgets on utility is similar, but overall, its effect is less pronounced compared to delay.
VI Conclusion
In this paper, we have studied the joint optimization of user association and resource allocation in multi-cell semantic communication networks. We introduced the concept of adaptive semantic communication and proposed a three-stage algorithm to maximize the total utility. Numerical results indicate that the performance of semantic communication can significantly outperform the traditional communication paradigm if we choose appropriate control decisions. In our future work, we will embark on building a real adaptive semantic communication system and refine our algorithm based on the experimental results in practical scenarios.
Appendix A Proof of Theorem 1
According to condition (i), when WD $n$ is specified, the corresponding application $a$ is also determined. For convenience, we use $u_{n}$ to denote $u^{a}_{0}$. Since its utility function is binary, its utility can only be 00 or $u_{n}$. When its utility is 00, we do not need to associate it with any BS. Therefore, without loss of generality, we can assume $c_{n}=c^{a}_{0}$, $d_{n}=d^{a}_{0}$, and WD $n$ takes utility $u_{n}$ if it is associated with some BS $m$. Then the objective function of problem (10) can be written as $\sum_{n\in\mathcal{N}}\sum_{m\in\mathcal{M}}x^{m}_{n}u_{n}$.
According to condition (ii), the energy constraint (10c) is redundant and we can set $f_{n}=f^{max}_{n}$ and $P_{n}=P^{max}_{n}$. Since both $f_{n}$ and $c_{n}$ are fixed, the computation time $T^{c}_{n}$ becomes a constant. Therefore, the delay constraint (10b) is equivalent to $T^{t}_{n}\leq T^{max}-T^{c}_{n}$, where the right-hand side is a constant. By using condition (iii), the delay constraint (10b) can be further transformed to $z^{m}_{n}\geq x^{m}_{n}C_{n}$, where $C_{n}$ is a constant independent of $m$. Combining with constraint (5) and (10g) we can obtain an equivalent constraint on $x^{m}_{n}$:
$\sum_{n\in\mathcal{N}}x^{m}_{n}\lceil C_{n}\rceil\leq K_{m},\quad\forall m\in\mathcal{M},$
[…]
Appendix B Proof of Theorem 2
We prove by contradiction. We first show that (14) must be satisfied. If $E^{c}_{n}+E^{t}_{n}<E^{max}_{n}$, then we can assign the extra energy to the data transmission process, i.e. let $P_{n}^{\prime}=P_{n}^{*}+(E^{max}_{n}-E^{c}_{n}-E^{t}_{n})/T^{t}_{n}$. Apparently, $(f^{*}_{n},P_{n}^{\prime})$ is also a feasible solution of problem (12) but produces larger $d_{n}$ than $(f^{*}_{n},P^{*}_{n})$, contradicting the fact that $(f^{*}_{n},P^{*}_{n})$ is optimal.
Next, we show that (13) must be satisfied. If $T^{c}_{n}+T^{t}_{n}<T^{max}$, then we can assign the extra time to the computation process, i.e. let the new computation time be $T^{c,\prime}_{n}=T^{max}-T^{t}_{n}$. To process $c_{n}$ workload in $T^{c,\prime}_{n}$, the corresponding CPU frequency is $f_{n}^{\prime}=f_{n}^{*}-\frac{(T^{max}-T^{t}_{n}-T^{c}_{n})c_{n}}{T^{c}_{n}(T^{max}-T^{t}_{n})}$. Since $f_{n}^{\prime}<f_{n}^{*}$, we can save an amount of $\Delta E^{c}_{n}=c_{n}\gamma_{n}[(f_{n}^{*})^{2}-(f_{n}^{\prime})^{2}]$ energy for the computation process. Then the saved energy can be assigned to the communication process to achieve a larger $d_{n}$, i.e. $P_{n}^{\prime}=P^{*}_{n}+\Delta E^{c}_{n}/T^{t}_{n}$. One can easily verify that $(f_{n}^{\prime},P_{n}^{\prime})$ is feasible and produces larger $d_{n}$ than $(f_{n}^{*},P_{n}^{*})$, which results in a contradiction.
Appendix C Proof of Theorem 3
We only need to prove the second derivative is non-positive. For brevity, we temporarily drop the subscript $n$ and superscript $m$ in $\hat{z}^{m}_{n}$. Let $(c^{*}(\hat{z}),d^{*}(\hat{z}))$ be the optimal computation and communication workload pair when the allocated number of RBs is $\hat{z}$. Since $u^{a}(c,d)$ is non-descreasing, we can safely assume $d^{*}(\hat{z})=d^{max}(c^{*}(\hat{z}),\hat{z})$. Then we have
$\frac{\mathop{}\!\mathrm{d}{u}^{*}_{n}(\hat{z})}{\mathop{}\!\mathrm{d}\hat{z}}$ $=\frac{\mathop{}\!\mathrm{d}u^{a}\left(c^{*}(\hat{z}),d^{*}(\hat{z})\right)}{\mathop{}\!\mathrm{d}\hat{z}}$
$=\frac{\partial u^{a}}{\partial c^{*}}\times\frac{\mathop{}\!\mathrm{d}c^{*}}{\mathop{}\!\mathrm{d}\hat{z}}+\frac{\partial u^{a}}{\partial d^{*}}\times\frac{\mathop{}\!\mathrm{d}d^{*}}{\mathop{}\!\mathrm{d}\hat{z}}$
$=\frac{\partial u^{a}}{\partial c^{*}}\times\frac{\mathop{}\!\mathrm{d}c^{*}}{\mathop{}\!\mathrm{d}\hat{z}}+\frac{\partial u^{a}}{\partial d^{*}}\times\left(\frac{\partial d^{*}}{\partial c^{*}}\times\frac{\mathop{}\!\mathrm{d}c^{*}}{\mathop{}\!\mathrm{d}\hat{z}}+\frac{\partial d^{*}}{\partial\hat{z}}\right)$
$=\frac{\mathop{}\!\mathrm{d}c^{*}}{\mathop{}\!\mathrm{d}\hat{z}}\left(\frac{\partial u^{a}}{\partial c^{*}}+\frac{\partial u^{a}}{\partial d^{*}}\times\frac{\partial d^{*}}{\partial{c^{*}}}\right)+\frac{\partial u^{a}}{\partial d^{*}}\times\frac{\partial d^{*}}{\partial\hat{z}}.$ (22)
According to the discussion in Section IV-A, after plugging $f^{*}_{n}(c_{n})$ into $d_{n}$, $d^{max}_{n}$ equals $z^{m}_{n}$ multiplies a function of $c_{n}$, denoted by $g(c_{n})$. Therefore, we can express $d^{*}(\hat{z})$ as
$d^{*}(\hat{z})=\hat{z}g(c^{*}(\hat{z})).$ (23)
Since $c^{*}(\hat{z})$ is optimal, we have
$\frac{\mathop{}\!\mathrm{d}u^{a}}{\mathop{}\!\mathrm{d}c^{*}}=\frac{\partial u^{a}}{\partial c^{*}}+\frac{\partial u^{a}}{\partial d^{*}}\times\frac{\partial d^{*}}{\partial c^{*}}=0.$ (24)
Substituting (23) and (24) into (22) yields $\frac{\mathop{}\!\mathrm{d}{u}^{*}_{n}(\hat{z})}{\mathop{}\!\mathrm{d}\hat{z}}=\frac{\partial u^{a}}{\partial d^{*}}\times g(c^{*}(\hat{z}))$. Then the second derivative of ${u}^{*}_{n}(\hat{z})$ is
$\frac{\mathop{}\!\mathrm{d}^{2}{u}^{*}_{n}(\hat{z})}{\mathop{}\!\mathrm{d}\hat{z}^{2}}=\frac{\partial u^{a}}{\partial d^{*}}\times\frac{\partial g}{\partial c^{*}}\times\frac{\partial c^{*}}{\partial\hat{z}}.$ (25)
Since $u^{a}(c,d)$ is non-decreasing, the first term in (25) is non-negative. For the second term, it is apparent that $d^{max}(c)$ is non-increasing with respect to $c$, hence $g(c^{*}(\hat{z}))$ is also non-increasing with respect to $c^{*}(\hat{z})$, so the second term is non-positive. Therefore, we only need to prove the third term is non-negative. Let $\alpha>1$ be an arbitrary constant. It is easy to verify that $(c^{*}(\hat{z}),\alpha d^{*}(\hat{z}))$ is achievable when the allocated RB grows to $\alpha\hat{z}$. The partial derivative of $u^{a}(c,d)$ at $(c^{*}(\hat{z}),\alpha d^{*}(\hat{z}))$ satisfies
$\frac{\mathop{}\!\mathrm{d}u^{a}}{\mathop{}\!\mathrm{d}c^{*}}\Big{|}_{(c^{*}(\hat{z}),\alpha d^{*}(\hat{z}))}$
$\quad=\frac{\partial u^{a}}{\partial c^{*}}\Big{|}_{(c^{*},\alpha d^{*})}+\frac{\partial u^{a}}{\partial d^{*}}\Big{|}_{(c^{*},\alpha d^{*})}\times\frac{\partial d^{*}}{\partial c^{*}}\Big{|}_{(c^{*},\alpha\hat{z})}$
$\quad\geq\frac{\partial u^{a}}{\partial c^{*}}\Big{|}_{(c^{*},d^{*})}+\frac{1}{\alpha}\frac{\partial u^{a}}{\partial d^{*}}\Big{|}_{(c^{*},d^{*})}\times\alpha\frac{\partial d^{*}}{\partial c^{*}}\Big{|}_{(c^{*},\hat{z})}$
$\quad=\frac{\partial u^{a}}{\partial c^{*}}\Big{|}_{(c^{*},d^{*})}+\frac{\partial u^{a}}{\partial d^{*}}\Big{|}_{(c^{*},d^{*})}\times\frac{\partial d^{*}}{\partial c^{*}}\Big{|}_{(c^{*},\hat{z})}=0$
where the inequality is based on Assumption 2, Assumption 3, $\frac{\partial d^{*}}{\partial c^{*}}\leq 0$, and the fact that $\frac{\partial d^{*}}{\partial c^{*}}$ is linear with respect to $\hat{z}$. Since the partial derivative is non-negative, we will obtain a larger utility if we increase the value of $c^{*}(\hat{z})$. Hence, when the allocated RB is $\alpha\hat{z}$, the optimal computation workload is larger than $c^{*}(\hat{z})$, i.e. $c^{*}(\alpha\hat{z})\geq c^{*}(\hat{z})$. Since the value of $\alpha$ is arbitrary, we can conclude that $c^{*}(\hat{z})$ is non-decreasing with respect to $\hat{z}$, which completes the proof.

### 09 Mar 2026 v3

@@
-Joint User Association and Resource Allocation for Multi-Cell Networks with Adaptive Semantic Communication
+Joint User Association and Resource Allocation for Adaptive Semantic Communication in 5G and Beyond Networks
 Abstract
-Semantic communication is a promising communication paradigm that utilizes Deep Neural Networks (DNNs) to extract the information relevant to downstream tasks, hence significantly reducing the amount of transmitted data.
-In current practice, the semantic communication transmitter for a specific task is typically pre-trained and shared by all users.
-However, due to user heterogeneity, it is desirable to use different transmitters according to the available computational and communication resources of users.
-In this paper, we first show that it is possible to dynamically adjust the computational and communication overhead of DNN-based transmitters, thereby achieving adaptive semantic communication.
-After that, we investigate the user association and resource allocation problem in a multi-cell network where users are equipped with adaptive semantic communication transmitters.
-To solve this problem, we decompose it into three subproblems involving the scheduling of each user, the resource allocation of each base station (BS), and the user association between users and BSs.
-Then we solve each problem progressively based on the solution of the previous subproblem.
-The final algorithm can obtain near-optimal solutions in polynomial time.
-Numerical results show that our algorithm outperforms benchmarks under various situations.
-Index Terms:
-Semantic communication, resource allocation, user association, multi-cell networks
+Semantic communication (SemCom) has emerged as a promising paradigm that leverages Deep Neural Networks (DNNs) to extract task-relevant information, thereby substantially reducing the volume of transmitted data.
+In existing implementations, the semantic transceiver is typically pre-trained for a specific task and uniformly adopted by all users.
+However, due to user heterogeneity in computational and communication capabilities, employing a single, fixed semantic transceiver may degrade the coding efficiency and transmission robustness.
+To address this issue, we first demonstrate the feasibility of dynamically adjusting the computational and communication overhead of DNN-based semantic transceivers, enabling a more flexible paradigm referred to as Adaptive Semantic Communication (ASC).
+Building on this concept, we formulate a joint user association and resource allocation problem for ASC in 5G and beyond networks, aiming to maximize overall system utility under energy and latency constraints.
+However, the problem is very challenging due to the inherent interdependencies among decision variables.
+To tackle this complexity, we decompose the original problem into three subproblems: (i) ASC scheme selection for each user, (ii) spectrum allocation at each Small-cell Base Station (SBS), and (iii) user association across SBSs.
+Each subproblem is solved sequentially based on the solutions of the preceding stages.
+The proposed algorithm efficiently yields near-optimal solutions with polynomial-time complexity.
+Simulation results demonstrate our approach outperforms existing baselines under various situations.
 I Introduction
-It has been over seventy years since the inception of Shannon’s information theory [1].
-During this period, traditional wireless communication technologies, aiming at the accurate transmission of bits, have made significant advancements and approached the Shannon limit.
-However, with the emergence of modern applications such as virtual reality and smart cities, the amount of data that needs to be transmitted continues to grow at an unprecedented pace.
-To meet the transmission demands with limited communication resources, a new communication paradigm called semantic communication has been proposed [2, 3, 4].
-With the help of artificial intelligence technologies, semantic communication only transmits the information relevant to downstream tasks (which is usually referred to as semantic information), hence leading to a significant reduction in data traffic.
-Although semantic communication can reduce the amount of transmitted data, extracting semantic information introduces additional computational overhead.
-In current practice, users performing the same task typically use the same semantic communication transmitter.
-However, due to the heterogeneity of users and networks, it is desirable to dynamically adjust the transmitter according to the available computational and communication resources.
-For example, since the transmitters are implemented by Deep Neural Networks (DNNs), we can decrease the model size of DNNs for users with limited computational capabilities.
-Similarly, for users with poor channel conditions, we can further reduce the amount of transmitted data by only transmitting a portion of the semantic information.
-By combining these two methods, we can obtain a range of semantic communication schemes with different computational overhead, communication overhead, and result accuracies.
-By selecting the optimal semantic communication scheme for each user, we can achieve adaptive semantic communication and significantly improve the system performance.
-In this paper, we consider the joint user association and resource allocation problem for multi-cell networks with adaptive semantic communication.
-However, due to user heterogeneity (e.g.
-variations in computational capabilities, maximum transmitting power, and energy budget) and network heterogeneity (e.g.
-variations in user distribution, base station coverage, and channel conditions), there does not exist a straightforward solution for this problem.
-To solve this problem, we need to address the following challenges.
-  • Firstly, both computation and communication processes consume energy and time.
-Therefore, there exists a trade-off between the computational workload and the communication workload that a user can handle.
-Since the result accuracy depends on both kinds of workloads, we need to quantify this trade-off in order to obtain the optimal semantic communication scheme.
-However, how to quantify this trade-off is not explicitly clear.
-  • Secondly, in addition to transmitting power and transmission time, the spectrum bandwidth also influences the amount of transmitted data, thereby further affecting the result accuracy.
-However, due to user heterogeneity, the impact of spectrum bandwidth on result accuracy varies from user to user and hard to quantify.
-As a result, there is no simple rule to find the optimal spectrum allocation.
-  • Thirdly, due to network heterogeneity, associating all users with the nearest base station (BS) may result in severe load imbalance.
-Therefore, to make appropriate user association decisions, we need to consider both load balancing and channel conditions comprehensively.
-However, it is not clear how to measure the impact of these two factors.
-To address the above challenges, this paper designs an efficient approximate algorithm to obtain near-optimal user association and resource allocation decisions in polynomial time.
-Our main contributions can be summarized as follows:
-  • We propose the concept of adaptive semantic communication and formulate the user association and resource allocation problem in a multi-cell network.
-We also show the formulated problem is strongly NP-hard.
-  • We progressively solve the problem by decomposing it into three subproblems involving the scheduling of each user, the resource allocation of each BS, and the user association between users and BSs.
-The solution to each subproblem relies on the properties of the optimal solution of the previous subproblem.
-The final algorithm can obtain near-optimal solutions in polynomial time.
-  • We conduct extensive simulations to evaluate the proposed algorithm under various settings.
-Numerical results demonstrate that our algorithm achieves significant improvement over benchmark algorithms.
+The rapid proliferation of data-intensive applications—such as virtual reality, autonomous systems, and smart cities—has resulted in an unprecedented surge in data demand [2, 1].
+Owing to the intrinsic limitations of communication resources (e.g., spectrum and time), such demand has already surpassed the transmission capacity of the conventional communication paradigm, which primarily emphasizes the accurate delivery of raw bits.
+To address this challenge, recent research has investigated innovative approaches that reduce the transmission volume through the integration of Artificial Intelligence (AI) techniques.
+In particular, a novel communication paradigm known as semantic communication (SemCom) has been proposed [4, 26, 12].
+By leveraging AI, SemCom focuses on transmitting only the information that is meaningful and relevant to downstream tasks—commonly referred to as semantic information—thereby enabling substantial reductions in data traffic without compromising task performance.
+Although SemCom effectively reduces the volume of transmitted data, it also introduces non-negligible computational overhead due to the extraction of semantic information.
+In current implementations, users performing the same downstream task typically share a common SemCom transceiver.
+However, due to the inherent heterogeneity in user devices and network conditions, it is desirable to dynamically adapt the transceiver based on the available computational and communication resources.
+For instance, as semantic transceivers are commonly realized using Deep Neural Networks (DNNs), we can actively reduce the DNN model size to mitigate computational demands for users with constrained computational capabilities.
+Similarly, for users experiencing poor channel conditions, the volume of transmitted data can be further decreased by selectively transmitting a subset of the semantic information.
+By jointly applying these two strategies, a spectrum of SemCom schemes can be constructed, each characterized by different levels of computational complexity, communication overhead, and inference accuracy.
+Selecting the most appropriate scheme for each user leads to the concept of Adaptive Semantic Communication (ASC), which significantly enhances the efficiency of both communication and computaional resources.
+To demonstrate the potential of ASC in future communication systems, this paper investigates the performance optimization of ASC in 5G and beyond networks, where Small-cell Base Stations (SBSs) are densely deployed to meet the high-quality transmission demands of users.
+The proposed optimization framework simultaneously addresses user association, resource allocation, and ASC scheme selection to maximize system utility under energy and latency constraints.
+However, this problem is inherently challenging due to the complex interdependencies among decision variables.
+The primary challenges include:
+  • Computation-communication trade-off: For each user, both extraction and transmission of semantic information consume time and energy.
+With limited time and energy budgets, a fundamental trade-off exists between the computational and communication workloads that a user can handle.
+To identify the optimal ASC scheme, it is crucial to characterize this trade-off and understand its effect on the inference accuracy of downstream tasks.
+  • Spectrum sharing among heterogeneous users: Users associated with the same SBS must share limited spectral resources.
+However, due to user heterogeneity—such as variations in channel conditions, utility functions, and computational capabilities—the marginal utility of spectrum differs across users, making optimal spectrum allocation a non-trivial problem.
+  • User association under spatial and network heterogeneity: In 5G and beyond networks, individual users are typically within the coverage of multiple SBSs.
+However, associating each user with its nearest SBS can lead to significant load imbalance due to the uneven spatial distribution of users and SBSs.
+Hence, optimizing user association requires a joint consideration of interrelated factors such as load balancing, channel conditions, and user-specific transmission demands.
+  • Coupled decision variables across network layers: The decisions variables at different network levels are also deeply coupled.
+For example, user association determines the grouping of users that share spectrum resources.
+Furthermore, the spectrum allocation impacts each user’s transmission capacity, which influences the trade-off between communication and computation, and ultimately affects the choice of ASC schemes.
+To address the aforementioned challenges, we decompose the original optimization problem into three hierarchical subproblems, each corresponding to a different network layer:
+  • ASC Scheme Selection: We first focus on selecting the optimal ASC scheme for each user under fixed user association and spectrum allocation.
+To characterize the computation-communication trade-off, we formally prove that both energy and latency constraints are tight at the optimal solution.
+This key property enables us to efficiently determine the maximum transmittable data given a fixed computational workload.
+Building upon this result, we can derive the optimal computational workload and subsequently determine the optimal ASC scheme.
+  • Spectrum Allocation: Next, we address the spectrum allocation under fixed user association.
+We consider two classes of utility functions: concave and general.
+For the concave case, we demonstrate that a simple greedy strategy yields the optimal allocation.
+For general utility functions, we develop a dynamic programming-based approach to efficiently determine the optimal solution.
+  • User Association: Finally, we optimize user association across SBSs.
+The proposed method iteratively updates each user’s association starting from a dummy initialization.
+To mitigate computational complexity, we employ a linear approximation to estimate the marginal utility change for each potential reassignment.
+Furthermore, we introduce a kurtosis-based heuristic to determine the processing order of users.
+These subproblems are inherently interrelated, with each depending on the optimal solutions of the preceding layers.
+By systematically leveraging the structural characteristics of each subproblem and their interconnections, we eventually obtain an efficient algorithm capable of producing near-optimal solutions with polynomial time complexity.
+The main contributions of this paper are summarized as follows:
+  • We introduce the concept of ASC and discuss how to implement it in practice.
+Building on this foundation, we formulate a performance optimization problem for ASC in 5G and beyond networks, aiming to maximize the aggregate user utility subject to energy and delay constraints.
+Through formal analysis, we prove that the formulated problem is strongly NP-hard, highlighting its computational intractability.
+  • To address the complexity of the original problem, we decompose it into three interrelated subproblems: ASC scheme selection, spectrum allocation, and user association.
+By identifying the mutual dependencies among the optimal solutions of these subproblems, we develop an efficient algorithm capable of achieving near-optimal solutions within polynomial time complexity.
+  • We conduct extensive simulations to compare the proposed algorithm with both the conventional communication paradigm and existing SemCom schemes.
+The numerical results demonstrate that our approach significantly outperforms these benchmarks under diverse system configurations.
 The rest of the paper is organized as follows.
@@
 II Related Work
-User association is an effective approach to balance the workload among multiple BSs, hence improve the overall system performance.
-In [5], the authors formulate the user association problem as an integer programming prolem and develop an efficient iterative algorithm by relaxing the integral constraints.
-However, the obtained solution allows users to associate with multiple BSs, which is usually not realistic in practice.
-The authors in [6] consider the user association problem in the context of massive multiple-input-multiple-output networks.
-The objective is to maximize the network utility by optimizing the fraction of time that a user associate with each BS.
-This problem is further extended to the case with cooperative BSs in [7].
-Although user association is well studied in the conventional communication paradigm [8, 9, 10], these methods are not applicable in semantic communication because we also need to consider the influence of users’ computational capabilities on their transmission demands.
-The concept of semantic communication has been proposed for a long time [11].
-Recently, with the advancement of artificial intelligence technology, there has been a growing interest in utilizing DNNs to implement semantic communication.
-The authors in [12] designed a semantic communication framework for text transmission.
-Their approach is extended to construct semantic communication systems for other types of sources (e.g.
-image[13], speech[14], and video[15]), which has shown significant improvements over the conventional communication paradigm.
-To improve the performance of semantic communication systems, the authors in [16, 17, 18, 19, 20, 21] investigate the optimal resource allocation in various situations.
-However, they did not take into account the influence of users’ computational capabilities.
-Besides, most of these works only consider a single BS, which may lead to suboptimal result in ultra-dense networks.
+II-A User Association in 5G and Beyond Networks
+To meet the continuous growth in data traffic, telecommunication operators have been compelled to deploy a large number of SBSs, giving rise to the concept of Ultra-Dense Networks (UDNs) and Heterogeneous Netoworks (HetNets) [16, 30, 9, 36].
+Given the limited communication resources available at each SBS, the design of effective user association strategies is crucial for balancing network load across SBSs and enhancing overall system performance [3, 21].
+In [5], the authors formulated a user association problem aimed at maximizing network utility by optimizing the fraction of time each user is associated with different SBSs.
+This work was later extended to cooperative SBS scenarios in [40].
+To address dynamic network conditions, a robust user association algorithm was proposed in [19] based on predicted future traffic fluctuations.
+The long-term user association problem was studied in [44], where a matching game-based algorithm was developed to efficiently maximize data rate and minimize the number of handovers.
+User association is also commonly studied in conjunction with resource allocation—such as power and bandwidth—to achieve improved system performance [45, 25, 13, 29].
+More recently, AI techniques have been introduced to solve the user association problem.
+Specifically, [24] proposed a decentralized multi-agent actor-critic framework to maximize energy efficiency in UDNs, while [37] applied a deep Q-network to learn optimal user association strategies from historical experience.
+In [18], a deep learning-based approach was developed to jointly optimize transmit power and user association.
+Although user association is extensively investigated in existing literature, these studies are largely confined to the conventional communication paradigm.
+In contrast, our work introduces the ASC framework, which requires joint consideration of communication and computational resources, as well as their impact on the inference accuracy of downstream tasks.
+This added complexity significantly differentiates our problem from traditional user association formulations, thereby necessitating the development of novel algorithmic approaches.
+II-B Resource Allocation for Semantic Communication
+The concept of SemCom was originally introduced decades ago [28], but practical implementation remained elusive until recent advances in AI technologies enabled new possibilities.
+In particular, the emergence of DNNs has spurred growing interest in leveraging AI to realize SemCom systems.
+In [35], the authors proposed a pioneering SemCom framework for text transmission, which was subsequently extended to other modalities, including images [14], speech [33], and video [15].
+These implementations demonstrated substantial performance gains compared to conventional communication paradigms.
+To further enhance SemCom systems, several studies have explored semantic-aware resource allocation.
+For example, [39] investigated spectral efficiency in the semantic domain and re-examined resource allocation from a semantic perspective.
+In [38], the authors introduced the concept of semantic entropy to quantify task-relevant semantic information.
+They also developed a novel Quality-of-Experience (QoE) model for semantic-aware resource optimization and proposd a deep Q-network-based solution.
+Similarly, [47] designed an improved deep reinforcement learning (DRL) algorithm to enable server coordination during training, thereby accelerating convergence to a near-global optimum.
+The authors in [34] develop semantic channel models with a new performance metric and systematically optimize user association and bandwidth allocation.
+The work in [10] introduces a SemCom–empowered AIGC framework that leverages a resource-aware workload trade-off scheme to improve latency and content quality in dynamic wireless environments.
+Recent research has also recognized that not all semantic information contributes equally to the inference accuracy of downstream tasks.
+As a result, prioritizing the transmission of high-impact semantic components has emerged as a promising strategy.
+For instance, [31] proposed a DRL-based algorithm using proximal policy optimization to jointly determine both resource allocation and the subset of semantic information to be transmitted.
+In [46], a dynamic resource allocation policy was designed using DRL to ensure that data with higher semantic value receives preferential access to limited communication resources.
+The study in [20] further explored the joint optimization of semantic information compression and resource block allocation to maximize task success probability.
+The aforementioned studies primarily focus on the communication process, aiming to adaptively adjust communication overhead based on users’ available communication resources.
+However, in practical systems, users also exhibit significant heterogeneity in computational capabilities.
+The ASC framework proposed in this paper addresses this issue by simultaneously accommodating variations in both computational and communication resources across users.
+Building on this framework, we investigate the performance optimization of ASC in 5G and beyond networks, which necessitates the joint optimization of semantic scheme selection, user association, and resource allocation.
 III System Model and Problem Formulation
-In this section, we first present the considered system and describe the concept of adaptive semantic communication.
+In this section, we first present the considered system and describe the concept of ASC.
 After that, we formulate an utility maximization problem and prove it is strongly NP-hard.
+The major notations used in this paper are summarized in Table I.
+TABLE I: Major Notations
+Notation Description
+$M$ Number of SBSs.
+NN Number of WDs.
+AA Number of applications.
+$x^{m}_{n}$ Whether WD $n$ is associated with SBS $m$.
+$z^{m}_{n}$ Number of SBS $m$’s RBs allocated to WD $n$.
+$c_{n},d_{n}$ Computation and communication workload of WD $n$.
+$f_{n}$ CPU frequency of WD $n$.
+$P_{n}$ Transmission power of WD $n$.
+$u^{a}(c,d)$ Utility function of application $a$.
+$h^{m}_{n}$ Channel gain between WD $n$ and SBS $m$.
+$W$ Spectrum bandwidth of each RB.
+$K_{m}$ Number of RBs available at SBS $m$.
+$I_{m}$ Average interference at SBS $m$.
+$r^{m}_{n}$ Transmission rate from WD $n$ to SBS $m$.
+$T^{c}_{n},T^{t}_{n}$ Time consumption due to computation and transmission.
+$E^{c}_{n},E^{t}_{n}$ Energy consumption due to computation and transmission.
 III-A Network Model
 Figure 1: A simple example of the considered system.
-As shown in Fig.
-1, we consider a multi-cell network that consists of $M$ BSs and $N$ wireless devices (WDs), where one WD may communicate with multiple BSs.
-The set of all BSs and WDs are denoted by $\mathcal{M}$ and $\mathcal{N}$, respectively.
-There are a set $\mathcal{A}$ of $A$ applications running in the system.
-Let $\mathcal{N}_{a}\subseteq\mathcal{N}$ be the set of WDs belonging to application $a\in\mathcal{A}$.
-Each WD in $\mathcal{N}_{a}$ has a mount of raw data to be transmitted to BSs for further processing.
-For example, the WDs may be surveillance cameras that send captured photos to BSs for security risk detection.
-Due to limited wireless resources, the WDs will first extract the small-sized semantic information from the raw data (e.g.
-feature maps of the original photos) and only send the extracted semantic information to BSs.
-In our considered model, a WD may be within the coverage of multiple BSs, so we need to determine which BS it communicates with.
-Specifically, the user association is denoted by the binary variable $x^{m}_{n}$, where $x^{m}_{n}=1$ if WD $n$ is associated with BS $m$, and $x^{m}_{n}=0$ otherwise.
-We assume a WD can communicate with at most one BS so $x^{m}_{n}$ should satisfy
+As illustrated in Fig.
+1, we consider a 5G and beyond network comprising $M$ SBSs and $N$ wireless devices (WDs), where each WD may reside within the coverage area of multiple SBSs.
+The sets of SBSs and WDs are denoted by $\mathcal{M}$ and $\mathcal{N}$, respectively.
+The WDs may belong to various applications [^1] In this paper, application and downstream task are used interchangeably., such as intelligent surveillance and smart transportation.
+Let $\mathcal{A}$ represent the set of $A$ applications deployed in the system, and $\mathcal{N}_{a}\subseteq\mathcal{N}$ denote the subset of WDs associated with application $a\in\mathcal{A}$.
+Each WD in $\mathcal{N}_{a}$ generates raw data that must be transmitted to the SBSs for further processing.
+For instance, WDs may consist of surveillance cameras that upload captured images for security threat detection.
+Given the constraints of limited wireless resources, WDs perform semantic extraction locally to reduce data volume.
+Specifically, instead of transmitting raw data, WDs extract compact semantic representations (e.g., feature maps of the original images) and transmit only this semantic information to the SBSs.
+In our model, each WD may reside within the coverage area of multiple SBSs, so we need to decide which SBS the WD should associate with for communication.
+The user association is represented by a binary variable $x^{m}_{n}$, where $x^{m}_{n}=1$ indicates that WD $n$ is associated with SBS $m$, and $x^{m}_{n}=0$ otherwise.
+We assume that each WD can be associated with at most one SBS.
+Accordingly, the user association variables must satisfy the following constraint:
 $\sum_{m\in\mathcal{M}}x^{m}_{n}\leq 1,\quad\forall n\in\mathcal{N}.$ (1)
 III-B Adaptive Semantic Communication Model
-In existing semantic communication works, it is commonly assumed that the same transmitter is shared by all WDs.
-However, due to user heterogeneity, different WDs may prefer different semantic communication schemes.
-For example, if a WD has weak computational power but good channel conditions and rich spectrum resources, it may transmit the raw data to the BS without extracting the semantic information.
-In contrast, if a WD has strong computational power but limited communication resources, it can fully process the raw data and only send the final computation results to the BS.
-Therefore, by choosing appropriate semantic communication schemes, we can optimize the utilization of communication and computational resources, thereby enhancing system performance.
-In practice, there are several ways to generate different semantic communication schemes.
-For example, the works in [22, 23, 24, 25] proposed various techniques to train DNNs of different model sizes in a single training process.
-Smaller models require less computation but also have worse accuracy.
-Since the semantic transmitters are implemented by DNNs, we can use these techniques to obtain a series of transmitters to achieve efficiency-accuracy trade-offs.
-Another method to derive multiple semantic communication schemes is through semantic compression [18, 19, 20, 21].
-The basic idea is, after extracting the semantic information, the transmitter evaluates the importance of each component of the extracted semantic information and only transmit those that contribute most to the result accuracy.
-The evaluation of semantic information induces extra computation but the amount of data to be transmitted can be significantly reduced.
-According to the above discussion, for a given application $a$, its result accuracy is determined by the computation workload $c^{a}$ (defined as the number of CPU cycles required to extract the semantic information) and the communication workload $d^{a}$ (defined as the amount of data to be transmitted).
-Therefore, we can use the utility function $u^{a}(c,d)$ to represent the obtained utility when the computation and communication workload is $c$ and $d$.
-It is worth noting that in practice, $c$ and $d$ are often limited to discrete values.
-For example, in semantic compression, each component of the semantic information cannot be further divided so $d$ must be an integer multiple of the data size of each component.
-However, the possible values of $c$ and $d$ are usually relatively dense so $u^{a}(c,d)$ can be regarded as a sufficiently good approximation.
-III-C Computation and Transmission Model
-In current practice, semantic information extraction is usually performed by DNNs, which requires a large amount of computation.
-For energy efficiency considerations, we assume WDs adopt the Dynamic Voltage and Frequency Scaling (DVFS) technique [26] to dynamically adjust the CPU frequency.
-For WD $n\in\mathcal{N}_{a}$, let $f_{n}$ be its CPU frequency and $c_{n}$ be its computation workload, then the computation time for semantic information extraction is
+In SemCom systems, the desired semantic information varies across different downstream tasks.
+As a result, a DNN-based semantic transceiver has to be trained for each specific downstream task.
+Once trained, the transceiver is deployed to all WDs performing the corresponding task.
+However, due to the inherent heterogeneity in user capabilities, different WDs may prefer transceivers with varying computational and communication demands.
+Consider two illustrative extreme cases: a WD with limited computational capacity but favorable channel conditions and abundant spectrum resources may find it preferable to transmit raw data directly to the SBS, bypassing local semantic extraction.
+In contrast, a WD with strong computational capability but constrained communication resources may choose to fully process the data locally and transmit only the final inference result.
+By selecting appropriate SemCom schemes tailored to device capabilities, the system can achieve more efficient utilization of both communication and computational resources, thereby improving overall performance.
+In practical implementations, various techniques can be employed to generate multiple SemCom schemes.
+For instance, methods proposed in [43, 41, 42, 7] enable training of DNNs with variable model sizes within a single training process.
+Smaller models reduce computational cost at the expense of lower inference accuracy.
+Given that semantic transceivers are implemented using DNNs, such techniques can be directly applied to produce a series of transceivers offering different trade-offs between efficiency and accuracy.
+An alternative approach to constructing diverse SemCom schemes involves semantic compression [31, 46, 47, 20].
+This technique involves evaluating the significance of individual components within the extracted semantic information and selectively transmitting only those that contribute most to the accuracy of downstream inference.
+Although semantic compression introduces additional computational overhead for evaluating component importance, it can substantially reduce the amount of data to be transmitted.
+Based on the above discussion, for a given application $a$, the resulting inference accuracy depends on the computation workload $c$ (defined as the number of CPU cycles required for semantic extraction) and the communication workload $d$ (defined as the volume of data to be transmitted).
+Accordingly, we define a utility function $u^{a}(c,d)$ to represent the utility derived from using a SemCom scheme with workloads $c$ and $d$.
+It is important to note that in practice, the values of $c$ and $d$ are typically discrete.
+For example, in semantic compression, the semantic information is composed of discrete components that cannot be further subdivided, implying that $d$ must be an integer multiple of the size of each component.
+Nevertheless, since the feasible values of $c$ and $d$ are often densely distributed, the utility function $u^{a}(c,d)$ can still serve as a sufficiently accurate approximation in optimization formulations.
+III-C Computation and Communication Model
+In current practice, semantic information extraction is typically performed using DNNs, which entail substantial computational workloads.
+To improve energy efficiency, we assume that WDs employ the dynamic voltage and frequency scaling technique [32], which enables dynamic adjustment of CPU frequency based on processing demands.
+For WD $n\in\mathcal{N}_{a}$, let $f_{n}$ be its CPU frequency and $c_{n}$ be its computational workload, measured in CPU cycles.
+Then the computation time required for semantic information extraction can be expressed as:
 $T^{c}_{n}=\frac{c_{n}}{f_{n}}.$ (2)
-According to the results in [27], the energy consumption per CPU cycle is proportional to the square of the frequency, which can be expressed as $\gamma_{n}f^{2}_{n}$, where $\gamma_{n}$ is the energy-efficiency factor determined by the chip architecture.
-Hence, the energy consumption due to computation is
+As established in [8], the energy consumption per CPU cycle is proportional to the square of the frequency. Accordingly, the total energy consumption for computation is given by:
 $E^{c}_{n}=c_{n}\gamma_{n}f^{2}_{n}.$ (3)
-The spectrum bandwidth of each BS is equally divided into resource blocks (RBs) with bandwidth $W$ and the number of available RBs at BS $m$ is $K_{m}$.
-Let $z^{m}_{n}\in\{0,1,\dots,K_{m}\}$ be the number of BS $m$’s RBs allocated to WD $n$.
-As in [28, 29, 30, 31], we assume the inter-BS interference is randomized and/or mitigated by various coordination techniques so the total interference seen by each BS can be regarded as static.
-According to the Shannon capacity formula, the transmission rate from WD $n$ to BS $m$ can be expressed as:
+where $\gamma_{n}$ is a hardware-dependent energy efficiency coefficient determined by the chip architecture.
+On the communication side, we assume that the available spectrum at each SBS is equally divided into resource blocks (RBs) with a fixed bandwidth $W$.
+The total number of RBs available at SBS $m$ is denoted by $K_{m}$.
+Let $z^{m}_{n}\in\{0,1,\dots,K_{m}\}$ represent the number of SBS $m$’s RBs allocated to WD $n$.
+In addition, the uplink interference at the SBS is typically the superposition of a large number of signals from spatially distributed users.
+Due to high user density and effective interference mitigation techniques (e.g., SIC, ZF, MMSE), the impact of any single user’s transmit power on the total interference is often negligible.
+Therefore, it is reasonable to regard the uplink interference from other cells as static [17, 27, 23, 11].
+Under this assumption, the achievable transmission rate from WD $n$ to SBS $m$ is given by the Shannon capacity formula:
 $r^{m}_{n}=z^{m}_{n}W\log_{2}\left(1+\frac{P_{n}h^{m}_{n}}{\sigma^{2}+I_{m}}\right),$ (4)
-where $P_{n}$ is the transmission power of WD $n$, $h^{m}_{n}$ is the channel gain between WD $n$ and BS $m$, $\sigma^{2}$ is the noise power at BS $m$, and $I_{m}$ is the average interference seen by BS $m$.
-Clearly, for each BS $m$, the number of totally allocated RBs cannot exceed $K_{m}$, so we have
+where $P_{n}$ is the transmission power of WD $n$, $h^{m}_{n}$ denotes the channel gain between WD $n$ and SBS $m$, $\sigma^{2}$ is the noise power at SBS $m$, and $I_{m}$ represents the average interference experienced by SBS $m$.
+To ensure resource feasibility, the total number of RBs allocated by each SBS must not exceed its capacity:
 $\sum_{n\in\mathcal{N}}z^{m}_{n}\leq K_{m},\quad\forall m\in\mathcal{M}.$ (5)
-For WD $n\in\mathcal{N}_{a}$, let $d_{n}$ be the amount of data to be transmitted, then the transmission time is
+For WD $n\in\mathcal{N}_{a}$, let $d_{n}$ be the amount of semantic data to be transmitted, then the transmission time can be calculated as:
 $T^{t}_{n}=\frac{d_{n}}{\sum_{m\in\mathcal{M}}x^{m}_{n}r^{m}_{n}}$ (6)
-and the energy consumption due to transmission is
+and the corresponding energy consumption is:
 $E^{t}_{n}=P_{n}T^{t}_{n}.$ (7)
 III-D Problem Formulation
-In this paper, we aim to maximize the total utility by jointly optimizing the user association and resource allocation under energy consumption, wireless bandwidth, and completion time constraints.
-Based on the models described above, the corresponding optimization problem can be formulated as follows:
+In this paper, we aim to maximize the overall system utility by jointly optimizing user association, resource allocation, and SemCom scheme, subject to constraints on energy consumption, spectrum availability, and delay.
+Based on the system models developed above, the corresponding optimization problem can be formulated as follows:
 $\max_{\begin{subarray}{c}\bm{x},\bm{z},\bm{f},\bm{P}\\ \bm{c},\bm{d}\end{subarray}}\quad$ $\sum_{a\in\mathcal{A}}\sum_{n\in\mathcal{N}_{a}}u^{a}(c_{n},d_{n})$ (8)
-[…]
+$s.t.\quad$ (1) and (5) (8a)
 $T^{c}_{n}+T^{t}_{n}\leq T^{max},\quad\forall n\in\mathcal{N}$ (8b)
@@
 $z^{m}_{n}\in\{0,1,\dots,K_{m}\},\quad\forall m\in\mathcal{M},\forall n\in\mathcal{N}$ (8g)
-where $\bm{x},\bm{z},\bm{f},\bm{P},\bm{c},\bm{d}$ are vetors of corresponding decision variables.
+Here, $\bm{x},\bm{z},\bm{f},\bm{P},\bm{c},\bm{d}$ are the vectors of decision variables.
 Constraint (8a) contains the user association and RB allocation constraints described above.
-Constraint (10b) ensures the semantic information transmission is completed before the maximum delay $T^{max}$.
-Constraint (10c) requires the total energy consumption of WD $n$ does not exceed a given threshold $E^{max}_{n}$.
-Constraints (10d), (10e), (10f), and (10g) define the ranges of feasible values for the decisions variables, where $f^{max}_{n}$ and $P^{max}_{n}$ are the maximum CPU frequency and transmission power of WD $n$, respectively.
-The formulated problem is Mixed-Integer Programming (MIP) and can be hard to solve.
-Specifically, we will show that our problem is strongly NP-hard, so it does not even have a pseudo-polynomial time algorithm.
-Before stating our result, we give the definition of binary utility functions: we say an utility function $u^{a}(c,d)$ is binary if $u^{a}(c,d)=u^{a}_{0}$ for all $(c,d)\geq(c^{a}_{0},d^{a}_{0})$ and $u^{a}(c,d)=0$ otherwise, where $u^{a}_{0},c^{a}_{0},d^{a}_{0}$ are positive constants.
+Constraint (8b) ensures that the total latency, comprising both computation and transmission time, does not exceed the maximum delay threshold $T^{max}$.
+Constraint (8c) limits the total energy consumption of each WD to a specified maximum $E^{max}_{n}$.
+Constraints (8d) and (8e) define the feasible ranges for CPU frequency and transmission power, bounded by their respective hardware limits $f^{max}_{n}$ and $P^{max}_{n}$.
+Finally, constraints (8f) and (8g) define the domain of binary user association variables and integer-valued RB allocation variables.
+The formulated optimization problem is a mixed-integer programming (MIP) problem, which is generally challenging to solve.
+In the literature, NP-hardness is commonly employed to characterize the inherent difficulty of such problems.
+However, establishing NP-hardness alone does not necessarily imply that a problem is intractable in practice.
+In particular, weakly NP-hard problems—such as the classical 0-1 knapsack problem—can be addressed using pseudo-polynomial time algorithms, and are often solved efficiently through dynamic programming [22].
+In contrast, we demonstrate that the problem under consideration is strongly NP-hard, which indicates it admits no pseudo-polynomial time algorithm and is fundamentally more complex to solve.
 Theorem 1.
-Problem (10) is strongly NP-hard even if (i) there is only one WD for each running application and its utility function is binary, (ii) the energy budget $E^{max}_{n}$ is sufficiently large, and (iii) the channel gain $h^{m}_{n}$ and interference $I_{m}$ for different BS $m$ is equal.
-Proof.
-Please see Appendix A. ∎
-Thereom 1 indicates that problem (10) is hard to solve even for a significantly simplified case.
-Therefore, in general, it is computationally intractable to obtain the exact solution of problem (10).
-In the next section, we will design an approximate algorithm that produces near-optimal solutions for problem (10) in polynomial time.
+Problem (8) is strongly NP-hard even if (i) there is only one WD for each running application, (ii) the energy budgets $E^{max}_{n}$ are sufficiently large, and (iii) the channel gains $h^{m}_{n}$ and interference levels $I_{m}$ are identical across all SBSs.
+Proof:
+We consider a specific class of two-valued utility functions, defined as $u^{a}(c,d)=u^{a}_{0}$ for all $(c,d)\geq(c^{a}_{0},d^{a}_{0})$ and $u^{a}(c,d)=0$ otherwise, where $u^{a}_{0},c^{a}_{0},d^{a}_{0}$ are application-specific positive constants.
+According to condition (i), when WD $n$ is specified, the corresponding application $a$ is also determined.
+For notational convenience, we omit the superscript $a$ and use $u_{n}$ to denote $u^{a}_{0}$.
+Since the utility function is two-valued, WD $n$’s utility is either 0 or $u_{n}$.
+If the utility is 0, there is no need to associate it with any SBS.
+Therefore, without loss of generality, we can assume $c_{n}=c^{a}_{0}$, $d_{n}=d^{a}_{0}$, and WD $n$ takes utility $u_{n}$ if it is associated with some SBS $m$.
+Then the objective function of problem (8) can be re-written as $\sum_{n\in\mathcal{N}}\sum_{m\in\mathcal{M}}x^{m}_{n}u_{n}$.
+According to condition (ii), the energy constraint (8c) is redundant, allowing us to fix the CPU frequency and transmission power at their maximum values, i.e., $f_{n}=f^{max}_{n}$ and $P_{n}=P^{max}_{n}$.
+Since both $f_{n}$ and $c_{n}$ are fixed, the computation time $T^{c}_{n}$ becomes a constant.
+Therefore, the delay constraint (8b) is equivalent to $T^{t}_{n}\leq T^{max}-T^{c}_{n}$, where the right-hand side is a constant.
+By substituting the expression for $T^{t}_{n}$ from (6), this constraint can be interpreted as a lower bound on the number of RBs allocated to WD $n$.
+Consequently, combining with condition (iii), the delay constraint (8b) can be equivalently expressed as: $z^{m}_{n}\geq x^{m}_{n}C_{n}$, where $C_{n}$ is a constant independent of $m$.
+Combining with constraints (5) and (8g), we derive an equivalent constraint on $x^{m}_{n}$:
+$\sum_{n\in\mathcal{N}}x^{m}_{n}\lceil C_{n}\rceil\leq K_{m},\quad\forall m\in\mathcal{M},$
+where $\lceil\cdot\rceil$ is the ceiling function that returns the smallest integer which is greater than or equal to the input value.
+As a result, the original problem (8) reduces to a classical 0-11 multiple knapsack problem, which is known to be strongly NP-hard [22].
+∎
+Theorem 1 highlights the intrinsic complexity of problem (8), showing that it remains computationally intractable even in highly simplified settings.
+Consequently, it is generally infeasible to obtain the exact optimal solution for non-trivial network scale.
+To address this issue, we propose an efficient algorithm capable of producing near-optimal solutions in polynomial time.
 IV Algorithm Design and Performance Analysis
-In this section, we will progressively solve the formulated problem through three stages.
-Firstly, we find the optimal scheduling decisions $\bm{f},\bm{P},\bm{c},\bm{d}$ when the user association and RB allocation are given.
-Based on this result, we can solve the optimal RB allocation under any user association.
-Finally, we can obtain the optimal user association by employing a relaxation-then-refinement approach.
+In this section, we solve the formulated optimization problem through a three-stage framework.
+First, we determine the optimal scheduling variables $\bm{f},\bm{P},\bm{c},\bm{d}$ assuming fixed user association and RB allocation.
+Building on this result, we can optimize the RB allocation for any given user association.
+Finally, we derive the optimal user association by applying a relaxation-and-refinement strategy.
 IV-A Scheduling Subproblem for Each WD
-If the user association $\bm{x}$ and RB allocation $\bm{z}$ are given, problem (10) can be decomposed into $N$ subproblems concerning the scheduling of each WD.
-Specifically, for any WD $n\in\mathcal{N}_{a}$ associated with BS $m$, its scheduling subproblem can be formulated as:
+Given fixed user association $\bm{x}$ and RB allocation $\bm{z}$, the original problem (8) can be decomposed into $N$ independent subproblems, each corresponding to the scheduling decision for a single WD.
+Specifically, for any WD $n\in\mathcal{N}_{a}$ associated with SBS $m$, its local scheduling subproblem can be formulated as:
 $\max_{{f_{n}},{P_{n}},c_{n},d_{n}}\quad$ $u^{a}(c_{n},d_{n})$ (9)
@@
 $0\leq f_{n}\leq f^{max}_{n},0\leq P_{n}\leq P^{max}_{n}$ (9c)
-We say a computation and communication workload pair $(c_{n},d_{n})$ is achievable if there exists $f_{n},P_{n}$ such that constraints (11a)-(11c) are satisfied.
-The set of all achievable workload pairs is referred to as the achievable region.
-Since $u^{a}(c_{n},d_{n})$ is non-decreasing, the optimal workload pair $(c^{*}_{n},d^{*}_{n})$ must lie on the boundary of the achievable region, i.e.
-there is no achievable $(c_{n},d_{n})$ such that $c_{n}=c^{*}_{n}$ and $d_{n}>d^{*}_{n}$.
-To obtain the achievable region, we only need to calculate the maximum achievable $d_{n}$ for any given $c_{n}$, denoted by $d^{max}_{n}(c_{n})$.
-The calculation of $d^{max}_{n}(c_{n})$ can be formulated as another optimization problem
+We define a pair $(c_{n},d_{n})$ as achievable if there exists $f_{n}$ and $P_{n}$ such that the tuple $(c_{n},d_{n},f_{n},P_{n})$ satisfies constraints (9a)-(9c).
+The set of all achievable workload pairs constitutes the achievable region.
+Based on our discussion in Section III-B, we can assume the utility function $u^{a}(c_{n},d_{n})$ is non-decreasing, hence the optimal pair $(c^{*}_{n},d^{*}_{n})$ must lie on the boundary of the achievable region, i.e.
+there is no achievable pair $(c_{n},d_{n})$ such that $c_{n}=c^{*}_{n}$ and $d_{n}>d^{*}_{n}$.
+To characterize this boundary, it is sufficient to compute the maximum achievable $d_{n}$ for any given $c_{n}$, denoted by $d^{max}_{n}(c_{n})$.
+This process leads to the following optimization problem:
 $\max_{f_{n},P_{n}}\quad$ $d_{n}$ (10)
 […]
-To solve problem (12), we first show that its optimal solution satisfies the following condition.
+To solve problem (10), we first demonstrate that its optimal solution satisfies the following tightness condition:
 Theorem 2.
@@
 $E^{c}_{n}+E^{t}_{n}$ $=E^{max}_{n}.$ (12)
-Proof.
-Please see Appendix B. ∎
-According to Theorem 2, we can replace constraint (11a) and (11b) with (13) and (14).
-Substituting (2) and (6) into (13) and recalling that WD $n$ is associated with BS $m$ (i.e.
+Proof:
+We prove by contradiction. We first show that (12) must be satisfied. If $E^{c}_{n}+E^{t}_{n}<E^{max}_{n}$, then we can assign the extra energy to the data transmission process, i.e., let
+$P_{n}^{\prime}=P_{n}^{*}+\frac{E^{max}_{n}-E^{c}_{n}-E^{t}_{n}}{T^{t}_{n}}.$
+Apparently, $(f^{*}_{n},P_{n}^{\prime})$ is also a feasible solution of problem (10) but produces larger $d_{n}$ than $(f^{*}_{n},P^{*}_{n})$, contradicting the fact that $(f^{*}_{n},P^{*}_{n})$ is optimal.
+Next, we show that (11) must be satisfied.
+If $T^{c}_{n}+T^{t}_{n}<T^{max}$, then we can assign the extra time to the computation process, i.e., let the new computation time be $T^{c,\prime}_{n}=T^{max}-T^{t}_{n}$.
+To process $c_{n}$ workload in $T^{c,\prime}_{n},$ the corresponding CPU frequency is
+$f_{n}^{\prime}=f_{n}^{*}-\frac{(T^{max}-T^{t}_{n}-T^{c}_{n})c_{n}}{T^{c}_{n}(T^{max}-T^{t}_{n})}.$
+Since $f_{n}^{\prime}<f_{n}^{*}$, we can save an amount of $\Delta E^{c}_{n}=c_{n}\gamma_{n}[(f_{n}^{*})^{2}-(f_{n}^{\prime})^{2}]$ energy for the computation process.
+Then the saved energy can be assigned to the communication process to achieve a larger $d_{n}$, i.e.
+$P_{n}^{\prime}=P^{*}_{n}+\Delta E^{c}_{n}/T^{t}_{n}$.
+One can easily verify that $(f_{n}^{\prime},P_{n}^{\prime})$ is feasible and produces larger $d_{n}$ than $(f_{n}^{*},P_{n}^{*})$, which results in a contradiction.
+∎
+With Theorem 2, we can replace the inequality constraints (9a) and (9b) in problem (10) with equalities (11) and (12).
+Substituting (2) and (6) into (11) and using the fact that WD $n$ is associated with SBS $m$ (i.e.
 $x^{m}_{n}=1$) result in
 $P_{n}=\left(2^{\frac{d_{n}f_{n}}{(f_{n}T^{max}-c_{n})z^{m}_{n}W}}-1\right)\frac{\sigma^{2}+I_{m}}{h^{m}_{n}}.$ (13)
-Combining (15) with (11c) we get the feasible region of $f_{n}$:
+Combining (13) with (9c), we can derive the feasible region of $f_{n}$:
 $\frac{z^{m}_{n}Wc_{n}\log_{2}\left(1+\frac{P_{n}^{max}h^{m}_{n}}{\sigma^{2}+I_{m}}\right)}{z^{m}_{n}WT^{max}\log_{2}\left(1+\frac{P_{n}^{max}h^{m}_{n}}{\sigma^{2}+I_{m}}\right)-d_{n}}\leq f_{n}\leq f_{n}^{max}.$ (14)
@@
 $\quad\log_{2}\left(1+\frac{h^{m}_{n}(E^{max}-c_{n}\gamma_{n}f^{2}_{n})}{(T^{max}-\frac{c_{n}}{f_{n}})(\sigma^{2}+I_{m})}\right).$ (15)
-Hence, problem (12) is equivalent to finding the optimal $f_{n}$ that maximizes (17) subject to (16), which can be solved by letting $\frac{\partial d_{n}}{\partial f_{n}}=0$ and comparing the values of $d_{n}$ achieved by every critical point in the feasible region.
-Clearly, the optimal $f_{n}$ is a function of $c_{n}$, denoted by $f^{*}_{n}(c_{n})$.
-According to (17), we can regard $d_{n}$ as a function of $c_{n}$ and $f_{n}$.
-Then we have $d^{max}_{n}(c_{n})=\max_{f_{n}}d_{n}(c_{n},f_{n})=d_{n}(c_{n},f^{*}_{n}(c_{n}))$.
-Unfortunately, the expression of $\frac{\partial d_{n}}{\partial f_{n}}=0$ is fairly complex so we are unable to derive an analytical expression for $f^{*}_{n}(c_{n})$.
-However, we can still calculate the value of $f^{*}_{n}(c_{n})$ through numerical methods.
-Once we know how to calculate $d^{max}_{n}(c_{n})$, problem (11) is equivalent to $\max_{c_{n}}u^{a}(c_{n},d^{max}_{n}(c_{n}))$. Similarly, the optimal $c_{n}$ can be obtained by letting
+Thus, solving problem (10) reduces to finding the optimal $f_{n}$ (denoted as $f_{n}^{*}$) that maximizes (15) within the feasible region defined in (14).
+This can be achieved by letting $\frac{\partial d_{n}}{\partial f_{n}}=0$ and comparing the values of $d_{n}$ at every critical point in the feasible region.
+Although the analytical form of the stationary point (where $\frac{\partial d_{n}}{\partial f_{n}}=0$) is intractable due to the complexity of the expression, the value of $f^{*}_{n}$ can be efficiently computed using numerical techniques.
+[^2] In our implementation, we employ the fzero function in MATLAB, which uses Brent’s method [6], a hybrid root-finding approach combining bisection, secant and inverse quadratic interpolation methods.
+Denoting the optimal CPU frequency for a given $c_{n}$ as $f^{*}_{n}(c_{n})$, then we can express $d^{max}_{n}(c_{n})=\max_{f_{n}}d_{n}(c_{n},f_{n})=d_{n}(c_{n},f^{*}_{n}(c_{n}))$, where we regard $d_{n}$ as a function of $c_{n}$ and $f_{n}$ according to (15).
+Recall that the optimal pair $(c^{*}_{n},d^{*}_{n})$ must lie on the boundary of the achievable region, hence the problem (9) is equivalent to $\max_{c_{n}}u^{a}(c_{n},d^{max}_{n}(c_{n}))$.
+Similar to $f^{*}_{n}$, the optimal $c^{*}_{n}$ can be obtained by numerically solving
 $\frac{\mathop{}\!\mathrm{d}u^{a}(c_{n},d^{max}_{n}(c_{n}))}{\mathop{}\!\mathrm{d}c_{n}}$
@@
 $\quad=\frac{\partial u^{a}}{\partial c_{n}}+\frac{\partial u^{a}}{\partial d^{max}_{n}}\times\left(\frac{\partial d_{n}}{\partial c_{n}}+\frac{\partial d_{n}}{\partial f^{*}_{n}}\times\frac{\mathop{}\!\mathrm{d}f^{*}_{n}}{\mathop{}\!\mathrm{d}c_{n}}\right)$
-$\quad=\frac{\partial u^{a}}{\partial c_{n}}+\frac{\partial u^{a}}{\partial d^{max}_{n}}\times\frac{\partial d_{n}}{\partial c_{n}}=0$
+$\quad=\frac{\partial u^{a}}{\partial c_{n}}+\frac{\partial u^{a}}{\partial d^{max}_{n}}\times\frac{\partial d_{n}}{\partial c_{n}}=0,$
 where the second equality holds because $d^{max}_{n}(c_{n})=d_{n}(c_{n},f^{*}_{n}(c_{n}))$ and the third equality holds because $\frac{\partial d_{n}}{\partial f^{*}_{n}}=0$.
-Similarly, the optimal $c_{n}$ can only be solved by numerical methods.
-After obtaining the optimal $c_{n}$, the optimal values of the rest decision variables can be calculated easily.
-IV-B RB Allocation Subproblem for Each BS
-In the previous subsection, we have obtained the optimal $\bm{f},\bm{P},\bm{c},\bm{d}$ for any given $\bm{x}$ and $\bm{z}$.
-In this subsection, we still assume $\bm{x}$ is given and show how to find the optimal $\bm{z}$.
-Let $u^{*}_{n}(z^{m}_{n})$ be the optimal objective value of problem (11).
-The meaning of $u^{*}_{n}(z^{m}_{n})$ is the maximum achievable utility of WD $n$ when $z^{m}_{n}$ RBs are allocated to it.
-Since $\bm{x}$ is given, we can use $\mathcal{N}_{m}$ to denote the set of WDs associated with BS $m$.
-Then the RB allocation subproblem for BS $m$ becomes
+Once $c^{*}_{n}$ is found, the remaining optimal scheduling decisions $f^{*}_{n},P^{*}_{n},d^{*}_{n}$ can be readily computed.
+IV-B RB Allocation Subproblem for Each SBS
+In the previous subsection, we derived the optimal scheduling decisions $\bm{f},\bm{P},\bm{c},\bm{d}$ under given user association $\bm{x}$ and RB allocation $\bm{z}$.
+In this subsection, we continue to assume that $\bm{x}$ is fixed and focus on determining the optimal RB allocation $\bm{z}$.
+Let $u^{*}_{n}(z^{m}_{n})$ denote the maximum utility achieved by WD $n$ (i.e., the optimal objective value of problem (9)) when it is allocated $z^{m}_{n}$ RBs.
+Given the user association $\bm{x}$, we define $\mathcal{N}_{m}=\{n\in\mathcal{N}\ |\ x^{m}_{n}=1\}$ as the set of WDs associated with SBS $m$.
+Then the RB allocation subproblem for SBS $m$ can be formulated as:
 $\max_{z^{m}_{n}}\quad$ $\sum_{n\in\mathcal{N}_{m}}u^{*}_{n}(z^{m}_{n})$ (16)
@@
 $z^{m}_{n}\in\{0,1,\dots,K_{m}\},\quad\forall n\in\mathcal{N}_{m}.$ (16b)
-In this subsection, we will design two different algorithms for problem (18) based on the properties of the utility function $u^{a}(c,d)$.
-In practice, the most commonly used utility function is the result accuracy, e.g.
-the classification accuracy in image recognition tasks.
-In this case, the utility function usually exhibits concavity and other properties that can be exploited to accelerate the solving of problem (18).
-However, in some scenarios, we may encounter more general utility functions.
-Therefore, we also design an algorithm that does not require the concavity of the utility function.
+In this subsection, we will present two algorithms for problem (16) tailored to different classes of utility function $u^{a}(c,d)$.
+In many practical applications, such as image recognition and object detection, the utility usually corresponds to inference accuracy, which typically exhibits concavity and other exploitable properties.
+However, in some scenarios, we may encounter more general utility functions, for which alternative strategies are required.
 IV-B1 RB Allocation with Concave Utility
-If the utility function is the result accuracy or a concave function of the result accuracy, then it should satisfy the following three properties.
-Firstly, according to the results in [22, 23, 24, 25, 18, 19, 20, 21], the marginal return of $c$ and $d$ is diminishing.
-Hence, we can assume:
+If the utility function represents the inference accuracy or a concave function of the inference accuracy, then it usually satisfies the following three properties.
+Firstly, according to the results in [43, 41, 42, 7, 31, 46, 47, 20], the marginal returns of both computational and communication workloads are diminishing.
+Hence, we have:
 Assumption 1.
@@
 Secondly, the marginal return of computation workload should be non-decreasing with respect to the amount of transmitted data.
-This is because the improvement in result accuracy due to increased computation needs to be propagated to the destination through data transmission.
-If the amount of transmitted data is limited, then the enhancement obtained at the destination will be discounted.
-An extreme example is the marginal return of $c$ is 00 when $d=0$, as $u^{a}(c,0)=0$ for all $c$.
+This is because any improvement in inference accuracy resulting from additional computation will be diminished if insufficient information is conveyed to the destination.
+An extreme example is the marginal return $\frac{\partial u^{a}}{\partial c}$ remains zero when $d=0$, as $u^{a}(c,0)=0$ for all $c$.
 Therefore, we have:
@@
 $\frac{\partial u^{a}}{\partial d}|_{(c,\alpha d)}\leq\frac{1}{\alpha}\frac{\partial u^{a}}{\partial d}|_{(c,d)}$ for any $\alpha\geq 1$.
-Assumption 3 requires $\frac{\partial u^{a}}{\partial d}$ decreases at least at a linear rate of $d$.
-A common function that satisfies this assumption is the logarithmic function.
-In practice, this assumption generally does not hold strictly, but it can be regarded as approximately valid within the range of values that we are interested in.
-With the above assumptions, we can proceed to demonstrate the concavity of $u^{*}_{n}(z^{m}_{n})$.
-However, $z^{m}_{n}$ is a discrete variable and the concavity for discrete functions is not well-defined.
-Therefore, we relax $z^{m}_{n}$ to a real-value variable $\hat{z}^{m}_{n}$ and prove that $u^{*}_{n}(\hat{z}^{m}_{n})$ is concave.
+This assumption requires $\frac{\partial u^{a}}{\partial d}$ decreases at least at a linear rate of $d$.
+A common function that satisfies this property is the logarithmic function.
+Although Assumption 3 may not hold strictly in all scenarios, it can be regarded as approximately valid within the range of values we are interested in.
+Under these assumptions, we can proceed to demonstrate the concavity of $u^{*}_{n}(z^{m}_{n})$.
+Since the concavity for integer-valued variables is not well-defined, we relax $z^{m}_{n}$ to a continuous variable $\hat{z}^{m}_{n}\in[0,K_{m}]$ and prove that $u^{*}_{n}(\hat{z}^{m}_{n})$ is concave.
 Theorem 3.
 The function ${u}^{*}_{n}(\hat{z}^{m}_{n})$ is concave.
-Proof.
-Please see Appendix C. ∎
-Due to the concavity of $u^{*}_{n}(\hat{z}^{m}_{n})$, to solve (18), we only need to greedily allocate RBs to the WD that contributes most to the increase of utility.
-The detailed steps are as follows.
-At the beginning, we initialize $z^{m}_{n}=0$ for all $n\in\mathcal{N}_{m}$.
-After that, for each WD $n\in\mathcal{N}_{m}$, we calculate the increase of utility if an extra RB is allocated to it, i.e.
+Proof:
+[…]
+$\frac{\mathop{}\!\mathrm{d}{u}^{*}_{n}(\hat{z})}{\mathop{}\!\mathrm{d}\hat{z}}$ $=\frac{\mathop{}\!\mathrm{d}u^{a}\left(c^{*}(\hat{z}),d^{*}(\hat{z})\right)}{\mathop{}\!\mathrm{d}\hat{z}}$
+$=\frac{\partial u^{a}}{\partial c^{*}}\times\frac{\mathop{}\!\mathrm{d}c^{*}}{\mathop{}\!\mathrm{d}\hat{z}}+\frac{\partial u^{a}}{\partial d^{*}}\times\frac{\mathop{}\!\mathrm{d}d^{*}}{\mathop{}\!\mathrm{d}\hat{z}}$
+[…]
+$=\frac{\mathop{}\!\mathrm{d}c^{*}}{\mathop{}\!\mathrm{d}\hat{z}}\left(\frac{\partial u^{a}}{\partial c^{*}}+\frac{\partial u^{a}}{\partial d^{*}}\times\frac{\partial d^{*}}{\partial{c^{*}}}\right)+\frac{\partial u^{a}}{\partial d^{*}}\times\frac{\partial d^{*}}{\partial\hat{z}}$.
+(17)
+[…]
+$).d^{*}(\hat{z})=\hat{z}g(c^{*}(\hat{z}$)). (18)
+[…]
+$\frac{\mathop{}\!\mathrm{d}u^{a}}{\mathop{}\!\mathrm{d}c^{*}}=\frac{\partial u^{a}}{\partial c^{*}}+\frac{\partial u^{a}}{\partial d^{*}}\times\frac{\partial d^{*}}{\partial c^{*}}=0.$ (19)
+[…]
+$\frac{\mathop{}\!\mathrm{d}^{2}{u}^{*}_{n}(\hat{z})}{\mathop{}\!\mathrm{d}\hat{z}^{2}}=\frac{\partial u^{a}}{\partial d^{*}}\times\frac{\partial g}{\partial c^{*}}\times\frac{\partial c^{*}}{\partial\hat{z}}$.
+(20)
+[…]
+$\frac{\mathop{}\!\mathrm{d}u^{a}}{\mathop{}\!\mathrm{d}c^{*}}\Big|_{(c^{*}(\hat{z}),\alpha d^{*}(\hat{z}))}$
+$\quad=\frac{\partial u^{a}}{\partial c^{*}}\Big|_{(c^{*},\alpha d^{*})}+\frac{\partial u^{a}}{\partial d^{*}}\Big|_{(c^{*},\alpha d^{*})}\times\frac{\partial d^{*}}{\partial c^{*}}\Big|_{(c^{*},\alpha\hat{z})}$
+$\quad\geq\frac{\partial u^{a}}{\partial c^{*}}\Big|_{(c^{*},d^{*})}+\frac{1}{\alpha}\frac{\partial u^{a}}{\partial d^{*}}\Big|_{(c^{*},d^{*})}\times\alpha\frac{\partial d^{*}}{\partial c^{*}}\Big|_{(c^{*},\hat{z})}$
+$\quad=\frac{\partial u^{a}}{\partial c^{*}}\Big|_{(c^{*},d^{*})}+\frac{\partial u^{a}}{\partial d^{*}}\Big|_{(c^{*},d^{*})}\times\frac{\partial d^{*}}{\partial c^{*}}\Big|_{(c^{*},\hat{z})}=0$
+[…]
+Due to the concavity of $u^{*}_{n}(\hat{z}^{m}_{n})$, we can solve (16) by greedily allocating RBs to the WD that contributes most to the overall utility.
+The procedure is as follows.
+Initially, we set $z^{m}_{n}=0$ for all $n\in\mathcal{N}_{m}$.
+At each iteration, we compute the marginal utility gain for allocating an additional RB to each WD, i.e.
 $\Delta u_{n}=u^{*}_{n}(z^{m}_{n}+1)-u^{*}_{n}(z^{m}_{n})$.
-If WD $n^{\prime}$ has the greatest increase of utility, i.e.
-$n^{\prime}=\operatorname*{arg\,max}_{n\in\mathcal{N}_{m}}\Delta u_{n}$, we allocate one extra RB to it and update $z^{m}_{n^{\prime}}=z^{m}_{n^{\prime}}+1$.
-This process continues until all of the $K_{m}$ RBs are allocated.
+We then identify the WD $n^{\prime}=\operatorname*{arg\,max}_{n\in\mathcal{N}_{m}}\Delta u_{n}$ that yields the highest utility improvement and allocate one extra RB to it, i.e., $z^{m}_{n^{\prime}}=z^{m}_{n^{\prime}}+1$.
+This process is repeated until all $K_{m}$ RBs are allocated.
 Our algorithm is summarized in Algorithm 1.
-Notice that we only modify the values of $z^{m}_{n}$ for one WD at each iteration.
-Therefore, for the loop in line $3$-$5$, we only need to calculate $\Delta u_{n}$ for all $n\in\mathcal{N}_{m}$ for the first time we entered the loop.
-After that, we only need to update $\Delta u_{n}$ for the WD whose $z^{m}_{n}$ is modified in the previous iteration.
-Apparently, the time complexity of the whole algorithm is $O(|\mathcal{N}_{m}|K_{m})$, where $|\mathcal{N}_{m}|$ is the number of elements in $\mathcal{N}_{m}$.
+Note that in each iteration, only one variable $z^{m}_{n}$ is updated.
+Therefore, except for the initial iteration, where all $\Delta u_{n}$ values must be computed, subsequent iterations only require updating $\Delta u_{n}$ for the previously selected WD.
+The overall time complexity of the algorithm is $O(|\mathcal{N}_{m}|K_{m})$, where $|\mathcal{N}_{m}|$ is the number of elements in $\mathcal{N}_{m}$.
 Algorithm 1 Algorithm for the RB Allocation Subproblem with Concave Utility
@@
 For general utility functions, we cannot leverage the concavity of $u^{*}_{n}(\hat{z}^{m}_{n})$ to facilitate algorithm design.
-However, we can still solve the RB allocation subproblem (18) through dynamic programming.
-Specifically, let $u_{m}(k,j)$ be the maximum total utility of BS $m$ if there are only $k$ available RBs and $j$ WDs in BS $m$, i.e.
-$u_{m}(k,j)$ is the optimal objective value of the following optimization problem:
+However, the RB allocation subproblem (16) can still be effectively solved using dynamic programming.
+Specifically, let $U_{m}(k,j)$ be the maximum aggregate utility attainable at SBS $m$ when allocating $k$ RBs among the first $j$ WDs, where $j\in\{1,\dots,|\mathcal{N}_{m}|\}$.
+Formally, $U_{m}(k,j)$ corresponds to the optimal objective value of the following optimization problem:
 $\max_{z^{m}_{n}}\quad$ $\sum_{n=1}^{j}u^{*}_{n}(z^{m}_{n})$ (21)
@@
 $z^{m}_{n}\in\{0,1,\dots,k\},\quad\forall n\in\{1,\dots,j\}.$
-Apparently, the optimal objective value of problem (18) equals $u_{m}(K_{m},|\mathcal{N}_{m}|)$.
-For convenience, we use $z_{m}(k,j)$ to denote the optimal RB allocation for the last WD (i.e.
-WD $j$) in (19).
+Apparently, solving the RB allocation subproblem (16) is equivalent to computing $U_{m}(K_{m},|\mathcal{N}_{m}|)$.
+For notational convenience, we use $z^{*}_{m}(k,j)$ to denote the number of RBs allocated to the $j$-th WD under the optimal solution to problem (21).
 Clearly, when there is only one WD in the system, we should allocate all RBs to it.
 Therefore, we have
-$u_{m}(k,1)=u^{*}_{1}(k)\mbox{ and }z_{m}(k,1)=k.$
-Otherwise, for any $j\in\{2,3,\dots,N\}$, the value of $u_{m}(k,j)$ can be calculated through
-$u_{m}(k,j)=\max_{z^{m}_{j}}\{u_{m}(k-z^{m}_{j},j-1)+u^{*}_{j}(z^{m}_{j})\}.$ (20)
-Correspondingly, $z_{m}(k,j)$ is the optimal $z^{m}_{j}$ that maximizes (20).
-Based on (20), we can finally obtain $u_{m}(K_{m},|\mathcal{N}_{m}|)$.
-Then the optimal $z^{m}_{n}$ can be obtained by backtracking the optimal allocation decisions recorded in $z_{m}(k,j)$.
-Specifically, we can let $z^{m}_{|\mathcal{N}_{m}|}=z_{m}(K_{m},|\mathcal{N}_{m}|)$ and $z^{m}_{j}=z_{m}(K_{m}-\sum_{n=j+1}^{|\mathcal{N}_{m}|}z^{m}_{n},j)$ for $j\in\{1,\dots,|\mathcal{N}_{m}|-1\}$.
+$U_{m}(k,1)=u^{*}_{1}(k)\mbox{ and }z^{*}_{m}(k,1)=k.$
+For the case with two WDs, we consider all feasible RB allocations $z^{m}_{2}$ to the second WD and assign the remaining $k-z^{m}_{2}$ RBs to the first WD. The optimal allocation is thus given by:
+$U_{m}(k,2)=\max_{z^{m}_{2}}\left\{U_{m}(k-z^{m}_{2},1)+u^{*}_{j}(z^{m}_{2})\right\}$
+$z^{*}_{m}(k,2)=\operatorname*{arg\,max}_{z^{m}_{2}}\left\{U_{m}(k-z^{m}_{2},1)+u^{*}_{j}(z^{m}_{2})\right\}.$
+This recursive formulation can be generalized to any $j\in\{2,3,\dots,N\}$. Specifically, the dynamic programming recursion is given by:
+$U_{m}(k,j)=\max_{z^{m}_{j}}\left\{U_{m}(k-z^{m}_{j},j-1)+u^{*}_{j}(z^{m}_{j})\right\}$ (22)
+$z^{*}_{m}(k,j)=\operatorname*{arg\,max}_{z^{m}_{j}}\left\{U_{m}(k-z^{m}_{j},j-1)+u^{*}_{j}(z^{m}_{j})\right\}.$ (23)
+Based on (22), we can finally obtain $U_{m}(K_{m},|\mathcal{N}_{m}|)$ in a bottom-up manner.
+After that, the corresponding optimal RB allocation $z^{m}_{n}$ can be reconstructed via backtracking from the recorded $z^{*}_{m}(k,j)$.
+Specifically, the backtracking procedure begins with $z^{m}_{|\mathcal{N}_{m}|}=z^{*}_{m}(K_{m},|\mathcal{N}_{m}|)$ and for each $j\in\{1,\dots,|\mathcal{N}_{m}|-1\}$, we have
+$z^{m}_{j}=z^{*}_{m}\left(K_{m}-\sum_{n=j+1}^{|\mathcal{N}_{m}|}z^{m}_{n},j\right).$
 The detailed steps are summarized in Algorithm 2.
-When calculating $u_{m}(k,j)$ using (20), we need to iterate over all possible values of $z^{m}_{j}$, which amount to a total of $k$ values.
-Then one can easily verify that the time complexity of Algorithm 2 is $O\left(|\mathcal{N}_{m}|K^{2}_{m}\right)$.
+In each dynamic programming step (22), we iterate over all possible values of $z^{m}_{j}$, which amount to a total of $k$ values.
+Therefore, the overall computational complexity of Algorithm 2 is $O\left(|\mathcal{N}_{m}|K^{2}_{m}\right)$.
 Algorithm 2 Algorithm for the RB Allocation Subproblem with General Utility
-1:  Initialization: $u_{m}(k,1)=u^{*}_{1}(k),\ z_{m}(k,1)=k,\ \forall k\in\{1,\dots,K_{m}\}$;
+1: Initialization: $U_{m}(k,1)=u^{*}_{1}(k),\ z^{*}_{m}(k,1)=k,\ \forall k\in\{1,\dots,K_{m}\}$;
 2:for $j\in\{2,3,\dots,|\mathcal{N}_{m}|\}$ do
 3:  for $k\in\{0,1,\dots,K_{m}\}$ do
-4:        Calculate $u_{m}(k,j)$ according to (20) and record the corresponding $z_{m}(k,j)$.
+4:   Calculate $U_{m}(k,j)$ and $z^{*}_{m}(k,j)$ according to (22) and (23).
 5:  end for
@@
 8:for $j=\{|\mathcal{N}_{m}|,|\mathcal{N}_{m}|-1,\dots,1\}$ do
-9:     Let $z^{m}_{j}=z_{m}(k,j)$;
+9:  Let $z^{m}_{j}=z^{*}_{m}(k,j)$;
 10:  Let $k=k-z^{m}_{j}$;
@@
 IV-C User Association Subproblem
-In this subsection, we aim to find the optimal user association decision $\bm{x}^{*}$.
-We first assume that each WD is associated with all BSs, i.e.
+In this subsection, we address the optimization of the user association decision $\bm{x}$.
+We begin by considering a relaxed setting where each WD is temporarily associated with all available SBSs, i.e.
 $\mathcal{N}_{m}=\mathcal{N}$ for all $m\in\mathcal{M}$.
-By using algorithms developed in the previous subsections, we can calculate the optimal RB allocation and scheduling decisions.
-The resulting total utility is an upper bound for the objective value of the original problem (10).
-After that, we progressively transition to a user association decision that satisfies constraint (1) through $N$ steps.
-Specifically, in each step, we select a WD that does not satisfy constraint (1) and associate it with exactly one BS.
-Our objective is to minimize the decrease in utility at each step.
-The remaining problem is: (i) how to select a WD at each step and (ii) how to choose the target BS for the selected WD.
-We first address the second problem.
-Suppose we have selected a WD $n$ that is associated with all BSs.
-For an arbitrary BS $m$, let $(z^{m}_{n^{\prime}})_{n^{\prime}\in\mathcal{N}_{m}}$ be the optimal RB allocation.
-If WD $n$ is no longer associated with BS $m$, the RBs it occupies will be released and allocated to other WDs.
-The new optimal RB allocatation is denoted by $(\bar{z}^{m}_{n^{\prime}})_{n^{\prime}\in\mathcal{N}_{m}\setminus\{n\}}$.
-Then the decrease in utility for BS $m$ is
-$\Delta_{m}=u^{*}_{n}(z^{m}_{n})-\sum_{n^{\prime}\in\mathcal{N}_{m}\setminus\{n\}}\left(u^{*}_{n^{\prime}}(\bar{z}^{m}_{n^{\prime}})-u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})\right).$
-Recall that our objective is to minimize the decrease in total utility.
-Therefore, we only need to associate WD $n$ with the BS $m$ that has the greatest $\Delta_{m}$.
-When BS $m$ is associated with multiple WDs, the difference between $z^{m}_{n^{\prime}}$ and $\bar{z}^{m}_{n^{\prime}}$ is relatively small.
-Therefore, we can use the derivative of $u^{*}_{n^{\prime}}$ at $z^{m}_{n^{\prime}}$ to estimate the difference between $u^{*}_{n^{\prime}}(\bar{z}^{m}_{n^{\prime}})$ and $u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})$, i.e.
-$u^{*}_{n^{\prime}}(\bar{z}^{m}_{n^{\prime}})-u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})\approx\frac{\mathop{}\!\mathrm{d}u^{*}_{n^{\prime}}}{\mathop{}\!\mathrm{d}z^{m}_{n^{\prime}}}\times(\bar{z}^{m}_{n^{\prime}}-z^{m}_{n^{\prime}})$.
-If $z^{m}_{n^{\prime}}$ is continuous, then according to the KKT conditions, the derivative $\frac{\mathop{}\!\mathrm{d}u^{*}_{n^{\prime}}}{\mathop{}\!\mathrm{d}z^{m}_{n^{\prime}}}$ is equal for all $n^{\prime}$.
-In our problem, although $z^{m}_{n^{\prime}}$ takes discrete values, the values of derivatives should be close so we can regard them as approximately equal.
-Therefore, we only need to calculate the derivative for one $n^{\prime}$.
-Since it is not very convenient to calculate the derivatives for discrete variables, we use the marginal gain $\delta_{m}=u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}}+1)-u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})$ to replace the role of derivatives.
-Then we have
-$\Delta_{m}$ $\approx u^{*}_{n}(z^{m}_{n})-\sum_{n^{\prime}\in\mathcal{N}_{m}\setminus\{n\}}\delta_{m}(\bar{z}^{m}_{n^{\prime}}-z^{m}_{n^{\prime}})$
-$=u^{*}_{n}(z^{m}_{n})-\delta_{m}z^{m}_{n}$ (21)
-where the equality holds because $\sum_{n^{\prime}\in\mathcal{N}_{m}\setminus\{n\}}\bar{z}^{m}_{n^{\prime}}=K_{m}$ and $\sum_{n^{\prime}\in\mathcal{N}_{m}\setminus\{n\}}{z}^{m}_{n^{\prime}}=K_{m}-z^{m}_{n}$.
-The next problem is how to select the appropriate WD at each step.
-As described above, for any given WD $n$, we choose its target BS based on $\Delta_{m}$.
-However, the value of $\Delta_{m}$ depends on the set of WDs associated with BS $m$, which is $\mathcal{N}_{m}$.
-As our algorithm progresses, $\mathcal{N}_{m}$ also changes.
-Therefore, the BS previously chosen for WD $n$ may no longer be the optimal one.
-We refer to this phenomenon as “regret”.
-Intuitively, if we process the WDs in an appropriate order, the resulting regret will be relatively less.
-As mentioned above, regret occurs when the values of $\Delta_{m}$ change so that the previously maximum $\Delta_{m}$ no longer be the maximum.
-Notice that the values of $\Delta_{m}$ change in a gradual manner as $\mathcal{N}_{m}$ varies (i.e.
-$\Delta_{m}$ will not change abruptly), hence we have the following two observations.
-Firstly, for any given WD $n$, if the $\Delta_{m}$ for one BS $m$ is significantly higher than the rest BSs, then it is highly probable that BS $m$ is still the optimal choice for WD $n$ even if $\mathcal{N}_{m}$ changes.
-Secondly, if WD $n$ is processed relatively late, then the subsequent changes in $\mathcal{N}_{m}$ are limited.
-This implies that the changes in the values of $\Delta_{m}$ are also limited, hence it is unlikely to result in regret.
-Based on the above observations, to minimize regret, we should defer the processing of WDs whose values of $\Delta_{m}$ are close.
-To measure the degree of difference among $\Delta_{m}$, we define the kurtosis of WD $n$ as the ratio of the maximum $\Delta_{m}$ to the sum of all $\Delta_{m}$, i.e.
+Under this relaxed association, we apply the algorithms developed in previous subsections to compute the optimal RB allocation and scheduling decisions.
+The resulting total utility serves as an upper bound for the objective value of the original problem (8).
+Subsequently, we iteratively refine the user association decision to satisfy the constraint (1) through a sequence of $N$ steps.
+In each step, we select one WD that violates the association constraint and assign it to exactly one SBS.
+Our goal in each iteration is to minimize the resulting decrease in total utility.
+This process naturally raises two questions: (i) how to select the WD to process at each step, and (ii) how to determine its optimal SBS association.
+We first address the second question.
+Suppose we have selected WD $n^{\prime}$ which is currently associated with all SBSs.
+For an arbitrary SBS $m$, let $(z^{m}_{n})_{n\in\mathcal{N}_{m}}$ be its optimal RB allocation.
+If WD $n^{\prime}$ is disassociated from SBS $m$, the RBs it previously occupied can be redistributed to other WDs in $\mathcal{N}_{m}\setminus\{n^{\prime}\}$.
+The new optimal RB allocation is denoted by $(\bar{z}^{m}_{n})_{n\in\mathcal{N}_{m}\setminus\{n^{\prime}\}}$.
+Then the resulting decrease in utility for SBS $m$ is
+$\Delta_{m}$ $=\sum_{n\in\mathcal{N}_{m}}u^{*}_{n}(z^{m}_{n})-\sum_{n\in\mathcal{N}_{m}\setminus\{n^{\prime}\}}u^{*}_{n}(\bar{z}^{m}_{n})$
+$=u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})-\sum_{n\in\mathcal{N}_{m}\setminus\{n^{\prime}\}}\left(u^{*}_{n}(\bar{z}^{m}_{n})-u^{*}_{n}(z^{m}_{n})\right).$
+Hence, if we eventually associate WD $n^{\prime}$ with SBS $m^{\prime}$ (which means WD $n^{\prime}$ will disassociate with the rest SBSs), then the decrease in total utility will be $\Delta^{t}=\sum_{m\in\mathcal{M}\setminus\{m^{\prime}\}}\Delta_{m}$.
+Recall that our objective is to minimize $\Delta^{t}$, thus the optimal $m^{\prime}=\operatorname*{arg\,min}_{m^{\prime}}\Delta^{t}=\operatorname*{arg\,max}_{m^{\prime}}\Delta_{m^{\prime}}$.
+Since the differences between $z^{m}_{n}$ and $\bar{z}^{m}_{n}$ are typically small, we can employ linear approximation to estimate the change in utility due to a variation in the RB allocation:
+$u^{*}_{n}(\bar{z}^{m}_{n})-u^{*}_{n}(z^{m}_{n})\approx\frac{\mathop{}\!\mathrm{d}u^{*}_{n}}{\mathop{}\!\mathrm{d}z^{m}_{n}}\times(\bar{z}^{m}_{n}-z^{m}_{n}).$
+Since $z^{m}_{n}$ is the optimal RB allocation, the Karush-Kuhn-Tucker (KKT) conditions imply that the derivative $\frac{\mathop{}\!\mathrm{d}u^{*}_{n}}{\mathop{}\!\mathrm{d}z^{m}_{n}}$ should be equal for all $n\in\mathcal{N}_{m}$.
+[^3] Strictly speaking, the KKT conditions apply only to continuous variables.
+While $z^{m}_{n}$ takes discrete values, the corresponding derivatives are typically approximately equal in practice.
+Therefore, it suffices to evaluate the derivative for a single WD.
+Since calculating derivatives for discrete variables is often inconvenient or ill-defined, we approximate the derivative using the marginal gain:
+$\delta_{m}=u^{*}_{n}(z^{m}_{n}+1)-u^{*}_{n}(z^{m}_{n}).$
+Consequently, the utility reduction $\Delta_{m}$ can be expressed as:
+$\Delta_{m}$ $\approx u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})-\sum_{n\in\mathcal{N}_{m}\setminus\{n^{\prime}\}}\delta_{m}(\bar{z}^{m}_{n}-z^{m}_{n})$
+$=u^{*}_{n^{\prime}}(z^{m}_{n^{\prime}})-\delta_{m}z^{m}_{n^{\prime}}$ (24)
+where the equality holds because $\sum_{n\in\mathcal{N}_{m}\setminus\{n^{\prime}\}}\bar{z}^{m}_{n}=K_{m}$ and $\sum_{n\in\mathcal{N}_{m}\setminus\{n^{\prime}\}}{z}^{m}_{n}=K_{m}-z^{m}_{n^{\prime}}$.
+The next problem is how to select the appropriate WD to process at each step.
+As described above, for any given WD $n$, we choose its target SBS based on $\Delta_{m}$.
+However, the value of $\Delta_{m}$ depends on the set of WDs associated with SBS $m$, which is $\mathcal{N}_{m}$.
+As WDs progressively assigned to specific SBSs, the set $\mathcal{N}_{m}$ evolves, potentially altering the values of $\Delta_{m}$.
+Consequently, a previously optimal SBS for WD $n$ may no longer be optimal later—a phenomenon we term regret.
+To mitigate regret, we make the following two observations.
+First, if the value of $\Delta_{m}$ for one SBS significantly exceeds that of others, it is likely that the SBS remains optimal even as $\mathcal{N}_{m}$ changes.
+Second, if WD $n$ is processed late in the sequence, then the subsequent changes in $\mathcal{N}_{m}$ are limited.
+This implies that the variation of $\Delta_{m}$ is also limited, thereby reducing the risk of regret.
+Based on these observations, we prioritize WDs for which the values of $\Delta_{m}$ differ most distinctly. To quantify this, we define the kurtosis of WD $n$ as:
 $\kappa_{n}=\frac{\max_{m}\Delta_{m}}{\sum_{m}\Delta_{m}}.$
-Clearly, if the values of $\Delta_{m}$ are close, then the value of $\kappa_{n}$ is relatively small.
-To defer the processing of WDs with small $\kappa_{n}$, we only need to pick the WD with the largest $\kappa_{n}$ in each step.
-The complete algorithm is summarized in Algorithm 3, where $\mathcal{N}_{r}$ denotes the set of WDs to be processed.
-Let $K=\max_{m}K_{m}$ be the maximum number of RBs in all BSs.
-Based on the result in the previous subsection, the time complexity of line $3$ is $O(MNK)$ or $O(MNK^{2})$, depending on whether the utility function is concave.
-Since the time complexity of the loop in line $4$-$9$ is only $O(NM)$, the time complexity of the whole algorithm is $O(MN^{2}K)$ for concave utility functions and $O(MN^{2}K^{2})$ for general utility functions.
+A lower $\kappa_{n}$ indicates that $\Delta_{m}$ values are more uniform across SBSs, implying higher uncertainty in SBS selection.
+Hence, in each step, we select the WD with the largest $\kappa_{n}$, thereby deferring those with higher risk of regret.
+The overall procedure is outlined in Algorithm 3, where $\mathcal{N}_{r}$ denotes the set of WDs to be processed.
+Let $K=\max_{m}K_{m}$ be the maximum number of RBs in all SBSs.
+Based on the result in the previous subsection, calculating the optimal RB allocation in line $3$ has a time complexity of $O(MNK)$ for concave utility functions and $O(MNK^{2})$ for general ones.
+The time complexity of user selection and update process in line 44-99 is $O(NM)$.
+Thus, the total time complexity of Algorithm 3 is $O(MN^{2}K)$ for concave utility functions and $O(MN^{2}K^{2})$ for general utility functions.
 Algorithm 3 Algorithm for the User Association Subproblem
@@
 V Numerical Results
-TABLE I: Simulation Parameters
+TABLE II: Simulation Parameters
 Parameter Value
-Number of RBs in each BS $K_{m}$ $\mathcal{U}\{25,50,75,100\}$
+Number of RBs in each SBS $K_{m}$ $\mathcal{U}\{10,20,30,40\}$
 Bandwidth of each RB $W$ $0.2$ MHz
 Delay requirement $T^{max}$ $10$ ms
-Energy budget of WDs $E^{max}_{n}$ $2$ mJ
+Energy budget of WDs $E^{max}_{n}$ 22 mJ
 Maximum CPU frequency $f^{max}_{n}$ $\mathcal{U}[1,3]$ GHz
@@
 Energy efficiency of WDs’ processors $\gamma_{n}$ $\mathcal{U}[10^{-28},10^{-27}]$
-Noise power at BSs $\sigma^{2}$ $10^{-13}$
-Average interference at BSs $I_{m}$ $\mathcal{U}[10^{-13},10^{-12}]$
-Figure 2: Utility of each BS under the concave utility function.
-Figure 3: Utility of each BS under the general utility function.
-Figure 2: Utility of each BS under the concave utility function.
-Figure 3: Utility of each BS under the general utility function.
-Figure 4: Total utility of different algorithms.
-Figure 5: Total utility under different numbers of BSs.
-Figure 6: Total utility under different numbers of WDs.
-Figure 5: Total utility under different numbers of BSs.
-Figure 6: Total utility under different numbers of WDs.
-Figure 7: Total utility under different delay requirements.
-Figure 8: Total utility under different energy budgets.
-In this section, extensive simulations are conducted to evaluate the performance of our algorithm.
-We consider a wireless network consisting of $M=5$ BSs and $N=30$ WDs that are randomly located in a $\times 500$m square area.
-For the channel model, we assume the pathloss is $128.1+37.6\log_{10}Dist$ dB and the shadowing factor is $6$ dB, where $Dist$ is the distance (in km) between WDs and BSs.
+Noise power at SBSs $\sigma^{2}$ $10^{-13}$
+Average interference at SBSs $I_{m}$ $\mathcal{U}[10^{-13},10^{-12}]$
+In this section, extensive simulations are conducted to evaluate the performance of our algorithm and numerical results show that our method significantly outperforms the conventional communication paradigm and existing semantic compression approach.
+V-A Simulation Setup
+We consider a 5G network consisting of $M=10$ SBSs and $N=100$ WDs that are randomly located in a $1000\times 1000m$ square area.
+For the channel model, we assume the pathloss is $128.1+37.6\log_{10}Dist$ dB and the shadowing factor is $6$ dB, where $Dist$ is the distance (in km) between WDs and SBSs.
 Our utility function is based on the empirical results about the relationship between result accuracy and computational or communication workload.
@@
 To evaluate our algorithm comprehensively, we consider a concave utility function $u^{a}(c,d)=A$ and a general non-concave function $u^{a}(c,d)=\frac{1}{1-A}$.
-The values of rest parameters are listed in Table I.
+The values of rest parameters are listed in Table II.
 To demonstrate the effectiveness of the proposed algorithm (labeled as “Prop”), we compare it with the following benchmarks.
@@
 Unless otherwise specified, each benchmark only differs from our algorithm in one stage.
-  • Traditional Communication (TC): the raw data is transmitted to the BS without processing.
-  • Fixed Semantic Communication (FSC): each WD adopts a fixed semantic communication scheme.
+  • Traditional Communication (TC): the raw data is transmitted to the SBS without processing.
+  • Fixed Semantic Communication (FSC): each WD adopts a fixed SemCom scheme.
+  • Semantic Compression (SC): fixed computation with selective transmission of semantic information.
   • Average RB Allocation (ARB): the RBs are allocated to each associated WD evenly.
-  • Nearest User Association (NUA): the WDs are associated with the nearest BSs.
-  • FSC+ARB+NUA (FAN): we use the above three algorithms to produce decisions in each stage respectively.
+  • Nearest User Association (NUA): the WDs are associated with the nearest SBSs.
 For TC, we assume the size of raw data satisfies $D_{raw}\in\mathcal{U}[0.4M,0.8M]$ bits.
-If WD $n$ successfully transmits the raw data under the delay and energy constraints, then the BS can use the maximum DNN model to process the raw data, which leads to result accuracy $\eta^{a}_{2}$.
-For FSC, the utilized fixed semantic communication scheme is $c=D^{a}/2$ and $d=D^{a}/2$.
-The data presented in this section are the average of ten repeated experiments with different random seeds.
-In Fig.
+If WD $n$ successfully transmits the raw data under the delay and energy constraints, then the SBS can use the maximum DNN model to process the raw data, which leads to result accuracy $\eta^{a}_{2}$.
+For FSC, the utilized fixed SemCom scheme is $c=C^{a}/2$ and $d=D^{a}/2$.
+In SC schemes [31, 46, 20], the complete semantic representation is first extracted, resulting in $c=C^{a}$.
+Subsequently, only the most critical semantic information is transmitted, subject to constraints on communication resources, latency, and energy consumption.
+To further validate the effectiveness of our design, we introduce two ablation baselines, namely ARB and NUA, which are employed to isolate and highlight the contributions of the proposed RB allocation and user association strategies, respectively.
+The data presented in this section are the average of $30$ repeated experiments with different random seeds.
+Figure 2: Locations of WDs and SBSs.
+Figure 3: Utility of each SBS under the concave utility function.
+V-B Performance Comparison under a Specific Scenario
+To elucidate the strategic differences among the considered algorithms, we first analyze system performance under a representative scenario.
+Fig.
+2 illustrates the randomly generated positions of WDs and SBSs, while Fig.
+3 presents the utility achieved by each SBS under the concave utility function.
+In our simulations, the available wireless resources are highly constrained relative to the volume of raw data.
+As a result, either none or only a limited fraction of WDs can successfully transmit their raw data to SBSs, leading to degraded system utility under TC.
+The adoption of SemCom alleviates this issue by substantially reducing the amount of data to be transmitted.
+Within this framework, SC outperforms FSC owing to its more adaptable transmission strategy.
+Nevertheless, SC still suffers from fixed computational requirements, resulting in excessive latency and energy consumption.
+In contrast, ASC achieves a notable performance gain.
+Due to the heterogeneity of WDs, equal allocation of RBs across users is inefficient.
+Consequently, although the ablation baseline ARB achieves higher utility for certain SBSs, the overall network utility remains inferior to that of the proposed approach.
+A similar observation applies to the NUA strategy, which tends to yield near-optimal performance in regions with sparsely distributed SBSs.
+However, in areas where SBSs are densely located, systematic user association becomes crucial.
+For instance, by associating some WDs in proximity to SBSs $1$ and $2$ with SBS $6$, the proposed method further enhances overall network utility.
+Figure 4: Total utility of different algorithms.
+V-C Performance under Different Utility Functions
+To provide a comprehensive comparison of the overall performance of the considered algorithms, Fig.
+4 depicts their performance under both concave and general utility functions.
+For the concave utility function, the performance trends of the algorithms are consistent with the case study presented in the previous subsection.
+Moreover, most algorithms exhibit comparable performance across the two utility functions, with the exception of TC.
+This divergence arises because, under the general utility function, utility is defined as the reciprocal of one minus accuracy, which increases sharply as accuracy approaches unity.
+Consequently, TC benefits from this property, as it achieves particularly high utility when a WD successfully transmits the raw data to the SBS.
+In both utility settings, the proposed algorithm consistently achieves the highest utility, thereby demonstrating its robustness and near-optimality across different classes of utility functions.
+Figure 5: Total utility under different numbers of SBSs.
+Figure 6: Total utility under different numbers of WDs.
+V-D Impact of Network Scale
+To evaluate the influence of network scale on algorithmic performance, Fig.
 5 and Fig.
-4, we show the utility of each BS under the concave and general utility functions, respectively.
-In both cases, our algorithm achieves the largest utility at all BSs, except for BS $5$, where NUA outperforms our algorithm.
-This is because there are more WDs around BS $5$ and they are all associated with BS $5$ in NUA, hence leading to a high utility.
-In our algorithm, however, we will associate part of these WDs with other BSs to achieve better load balance.
-As a result, the total utility of all BSs (which is shown in Fig.
-4) produced by our algorithm is higher than that of NUA.
-Another algorithm that has comparable performance is ARB.
-This is not surprising because the marginal gain of result accuracy drops very quickly with the increase of $d$.
-Consequently, the marginal gain of utility also drops rapidly with the growth of allocated RBs.
-This guarantees the average allocation of RBs can achieve close-to-optimal results.
-Compared to the previous algorithms, the performance of TC and FSC degrades significantly.
-The problem in TC and FSC is similar: they both adopt fixed transmission schemes so they are unable to efficiently utilize resources when dealing with heterogeneous users and networks.
-An interesting observation is that the performance of FAN is even significantly worse than TC.
-This indicates that adopting semantic communication alone does not necessarily enhance system performance if the resource allocation is not optimized.
-According to Fig.
-4, most algorithms exhibit similar performance under the two different utility functions, except for TC.
-This is because, under the general utility function, the utility is the reciprocal of one minus accuracy, hence the utility increases very quickly when the accuracy approaches one.
-This is beneficial for TC as it achieves particularly high accuracy when a WD successfully transmits the raw data to the BS.
-To understand the impact of network scale on algorithm performance, Fig.
-7 and Fig.
-7 demonstrate the total utility of all algorithms under different numbers of BSs and WDs.
-Due to space limitations, we only present the data under the concave utility function, but the result is similar under the general utility function.
-From Fig.
-7, one can observe that the utility of all algorithms increases with the number of BSs.
-However, when the number of BSs exceeds 6, the utility gains for Prop, ARB, and NUA are very limited.
-In contrast, the algorithms that adopt fixed transmission schemes maintain a relatively high growth rate.
-However, in Fig.
-7, the growth trend is exactly the opposite.
-Specifically, the total utility of Prop, ARB, and NUA grows continuously, indicating their ability to serve more WDs with existing communication resources.
-The only algorithm showing a decreased utility is FAN.
-This is because FAN adopts the average allocation of RBs, so the number of RBs allocated to each WD becomes less when there are more WDs.
-Since FAN also employs a fixed transmission scheme, some WDs are unable to complete the transmission under increasingly less RBs, hence resulting in zero utility.
-In addition to network scale, we also investigated the impact of delay and energy constraints on the algorithm results, as shown in Fig.
+6 present the total utility of all algorithms under varying numbers of SBSs and WDs.
+For clarity, only results under the concave utility function are shown, as similar trends are observed for the general utility function.
+As illustrated in Fig.
+5, increasing the number of SBSs reduces the average distance between WDs and SBSs, thereby improving channel gains.
+Moreover, a larger number of SBSs also provides additional RBs.
+Consequently, the utility achieved by all algorithms increases with the number of SBSs.
+However, the performance improvement of TC remains marginal, highlighting its inefficiency.
+Among the ASC-based schemes, ARB exhibits a smaller performance gain compared with the proposed algorithm and NUA, primarily because its average RB allocation strategy fails to fully exploit the additional RBs, thereby diminishing the benefits of network densification.
+A similar trend is observed when varying the number of WDs.
+Given the limited wireless resources, systematic resource allocation becomes critical to achieving higher utility.
+Nonetheless, since both TC and FSC employ fixed transmission strategies, they lack enough flexibility in resource allocation, resulting in minimal performance gains.
+For the same reason, although SC adopts an adaptive transmission strategy, its fixed computation mechanism restricts flexibility compared to ASC paradigms.
+Consequently, its performance improvement is also limited relative to the proposed algorithm, NUA, and ARB.
+Figure 7: Total utility under different delay requirements.
+Figure 8: Total utility under different energy budgets.
+V-E Impact of Delay and Energy Constraints
+In addition to network scale, we further examine the effects of delay and energy constraints on algorithmic performance, as depicted in Fig.
 7 and Fig.
 8.
-The algorithms most sensitive to delay are FSC and FAN, because an increase in delay allows WDs to process data in a more energy-efficient manner, thereby improving the probability of successful transmission.
-The impact of energy budgets on utility is similar, but overall, its effect is less pronounced compared to delay.
+Among all schemes, SC exhibits the highest sensitivity to both constraints.
+This behavior arises because SC employs a fixed computation process that consumes a constant amount of time and energy; thus, any additional delay or energy budget can be entirely utilized for transmitting more semantic information, resulting in substantial marginal gains.
+Furthermore, the total utility of all algorithms is observed to be more sensitive to delay requirements than to energy budgets.
+This outcome is not surprising, as the transmitted data volume scales linearly with transmission time but only logarithmically with transmission power.
+Notice that our algorithm constantly outperforms all benchmarks in all cases, demonstrating its effectiveness under diverse scenarios.
 VI Conclusion
-In this paper, we have studied the joint optimization of user association and resource allocation in multi-cell semantic communication networks.
-We introduced the concept of adaptive semantic communication and proposed a three-stage algorithm to maximize the total utility.
-Numerical results indicate that the performance of semantic communication can significantly outperform the traditional communication paradigm if we choose appropriate control decisions.
-In our future work, we will embark on building a real adaptive semantic communication system and refine our algorithm based on the experimental results in practical scenarios.
-Appendix A Proof of Theorem 1
-According to condition (i), when WD $n$ is specified, the corresponding application $a$ is also determined.
-For convenience, we use $u_{n}$ to denote $u^{a}_{0}$.
-Since its utility function is binary, its utility can only be 00 or $u_{n}$.
-When its utility is 00, we do not need to associate it with any BS.
-Therefore, without loss of generality, we can assume $c_{n}=c^{a}_{0}$, $d_{n}=d^{a}_{0}$, and WD $n$ takes utility $u_{n}$ if it is associated with some BS $m$.
-Then the objective function of problem (10) can be written as $\sum_{n\in\mathcal{N}}\sum_{m\in\mathcal{M}}x^{m}_{n}u_{n}$.
-According to condition (ii), the energy constraint (10c) is redundant and we can set $f_{n}=f^{max}_{n}$ and $P_{n}=P^{max}_{n}$.
-Since both $f_{n}$ and $c_{n}$ are fixed, the computation time $T^{c}_{n}$ becomes a constant.
-Therefore, the delay constraint (10b) is equivalent to $T^{t}_{n}\leq T^{max}-T^{c}_{n}$, where the right-hand side is a constant.
-By using condition (iii), the delay constraint (10b) can be further transformed to $z^{m}_{n}\geq x^{m}_{n}C_{n}$, where $C_{n}$ is a constant independent of $m$.
-Combining with constraint (5) and (10g) we can obtain an equivalent constraint on $x^{m}_{n}$:
-$\sum_{n\in\mathcal{N}}x^{m}_{n}\lceil C_{n}\rceil\leq K_{m},\quad\forall m\in\mathcal{M},$
-[…]
-Appendix B Proof of Theorem 2
-We prove by contradiction.
-We first show that (14) must be satisfied.
-If $E^{c}_{n}+E^{t}_{n}<E^{max}_{n}$, then we can assign the extra energy to the data transmission process, i.e.
-let $P_{n}^{\prime}=P_{n}^{*}+(E^{max}_{n}-E^{c}_{n}-E^{t}_{n})/T^{t}_{n}$.
-Apparently, $(f^{*}_{n},P_{n}^{\prime})$ is also a feasible solution of problem (12) but produces larger $d_{n}$ than $(f^{*}_{n},P^{*}_{n})$, contradicting the fact that $(f^{*}_{n},P^{*}_{n})$ is optimal.
-Next, we show that (13) must be satisfied.
-If $T^{c}_{n}+T^{t}_{n}<T^{max}$, then we can assign the extra time to the computation process, i.e.
-let the new computation time be $T^{c,\prime}_{n}=T^{max}-T^{t}_{n}$.
-To process $c_{n}$ workload in $T^{c,\prime}_{n}$, the corresponding CPU frequency is $f_{n}^{\prime}=f_{n}^{*}-\frac{(T^{max}-T^{t}_{n}-T^{c}_{n})c_{n}}{T^{c}_{n}(T^{max}-T^{t}_{n})}$.
-Since $f_{n}^{\prime}<f_{n}^{*}$, we can save an amount of $\Delta E^{c}_{n}=c_{n}\gamma_{n}[(f_{n}^{*})^{2}-(f_{n}^{\prime})^{2}]$ energy for the computation process.
-Then the saved energy can be assigned to the communication process to achieve a larger $d_{n}$, i.e.
-$P_{n}^{\prime}=P^{*}_{n}+\Delta E^{c}_{n}/T^{t}_{n}$.
-One can easily verify that $(f_{n}^{\prime},P_{n}^{\prime})$ is feasible and produces larger $d_{n}$ than $(f_{n}^{*},P_{n}^{*})$, which results in a contradiction.
-Appendix C Proof of Theorem 3
-We only need to prove the second derivative is non-positive.
-For brevity, we temporarily drop the subscript $n$ and superscript $m$ in $\hat{z}^{m}_{n}$.
-Let $(c^{*}(\hat{z}),d^{*}(\hat{z}))$ be the optimal computation and communication workload pair when the allocated number of RBs is $\hat{z}$.
-Since $u^{a}(c,d)$ is non-descreasing, we can safely assume $d^{*}(\hat{z})=d^{max}(c^{*}(\hat{z}),\hat{z})$.
-Then we have
-$\frac{\mathop{}\!\mathrm{d}{u}^{*}_{n}(\hat{z})}{\mathop{}\!\mathrm{d}\hat{z}}$ $=\frac{\mathop{}\!\mathrm{d}u^{a}\left(c^{*}(\hat{z}),d^{*}(\hat{z})\right)}{\mathop{}\!\mathrm{d}\hat{z}}$
-$=\frac{\partial u^{a}}{\partial c^{*}}\times\frac{\mathop{}\!\mathrm{d}c^{*}}{\mathop{}\!\mathrm{d}\hat{z}}+\frac{\partial u^{a}}{\partial d^{*}}\times\frac{\mathop{}\!\mathrm{d}d^{*}}{\mathop{}\!\mathrm{d}\hat{z}}$
-$=\frac{\partial u^{a}}{\partial c^{*}}\times\frac{\mathop{}\!\mathrm{d}c^{*}}{\mathop{}\!\mathrm{d}\hat{z}}+\frac{\partial u^{a}}{\partial d^{*}}\times\left(\frac{\partial d^{*}}{\partial c^{*}}\times\frac{\mathop{}\!\mathrm{d}c^{*}}{\mathop{}\!\mathrm{d}\hat{z}}+\frac{\partial d^{*}}{\partial\hat{z}}\right)$
-$=\frac{\mathop{}\!\mathrm{d}c^{*}}{\mathop{}\!\mathrm{d}\hat{z}}\left(\frac{\partial u^{a}}{\partial c^{*}}+\frac{\partial u^{a}}{\partial d^{*}}\times\frac{\partial d^{*}}{\partial{c^{*}}}\right)+\frac{\partial u^{a}}{\partial d^{*}}\times\frac{\partial d^{*}}{\partial\hat{z}}.$ (22)
-According to the discussion in Section IV-A, after plugging $f^{*}_{n}(c_{n})$ into $d_{n}$, $d^{max}_{n}$ equals $z^{m}_{n}$ multiplies a function of $c_{n}$, denoted by $g(c_{n})$.
-Therefore, we can express $d^{*}(\hat{z})$ as
-$d^{*}(\hat{z})=\hat{z}g(c^{*}(\hat{z})).$ (23)
-Since $c^{*}(\hat{z})$ is optimal, we have
-$\frac{\mathop{}\!\mathrm{d}u^{a}}{\mathop{}\!\mathrm{d}c^{*}}=\frac{\partial u^{a}}{\partial c^{*}}+\frac{\partial u^{a}}{\partial d^{*}}\times\frac{\partial d^{*}}{\partial c^{*}}=0.$ (24)
-Substituting (23) and (24) into (22) yields $\frac{\mathop{}\!\mathrm{d}{u}^{*}_{n}(\hat{z})}{\mathop{}\!\mathrm{d}\hat{z}}=\frac{\partial u^{a}}{\partial d^{*}}\times g(c^{*}(\hat{z}))$.
-Then the second derivative of ${u}^{*}_{n}(\hat{z})$ is
-$\frac{\mathop{}\!\mathrm{d}^{2}{u}^{*}_{n}(\hat{z})}{\mathop{}\!\mathrm{d}\hat{z}^{2}}=\frac{\partial u^{a}}{\partial d^{*}}\times\frac{\partial g}{\partial c^{*}}\times\frac{\partial c^{*}}{\partial\hat{z}}.$ (25)
-Since $u^{a}(c,d)$ is non-decreasing, the first term in (25) is non-negative.
-For the second term, it is apparent that $d^{max}(c)$ is non-increasing with respect to $c$, hence $g(c^{*}(\hat{z}))$ is also non-increasing with respect to $c^{*}(\hat{z})$, so the second term is non-positive.
-Therefore, we only need to prove the third term is non-negative.
-Let $\alpha>1$ be an arbitrary constant.
-It is easy to verify that $(c^{*}(\hat{z}),\alpha d^{*}(\hat{z}))$ is achievable when the allocated RB grows to $\alpha\hat{z}$.
-The partial derivative of $u^{a}(c,d)$ at $(c^{*}(\hat{z}),\alpha d^{*}(\hat{z}))$ satisfies
-$\frac{\mathop{}\!\mathrm{d}u^{a}}{\mathop{}\!\mathrm{d}c^{*}}\Big{|}_{(c^{*}(\hat{z}),\alpha d^{*}(\hat{z}))}$
-$\quad=\frac{\partial u^{a}}{\partial c^{*}}\Big{|}_{(c^{*},\alpha d^{*})}+\frac{\partial u^{a}}{\partial d^{*}}\Big{|}_{(c^{*},\alpha d^{*})}\times\frac{\partial d^{*}}{\partial c^{*}}\Big{|}_{(c^{*},\alpha\hat{z})}$
-$\quad\geq\frac{\partial u^{a}}{\partial c^{*}}\Big{|}_{(c^{*},d^{*})}+\frac{1}{\alpha}\frac{\partial u^{a}}{\partial d^{*}}\Big{|}_{(c^{*},d^{*})}\times\alpha\frac{\partial d^{*}}{\partial c^{*}}\Big{|}_{(c^{*},\hat{z})}$
-$\quad=\frac{\partial u^{a}}{\partial c^{*}}\Big{|}_{(c^{*},d^{*})}+\frac{\partial u^{a}}{\partial d^{*}}\Big{|}_{(c^{*},d^{*})}\times\frac{\partial d^{*}}{\partial c^{*}}\Big{|}_{(c^{*},\hat{z})}=0$
-where the inequality is based on Assumption 2, Assumption 3, $\frac{\partial d^{*}}{\partial c^{*}}\leq 0$, and the fact that $\frac{\partial d^{*}}{\partial c^{*}}$ is linear with respect to $\hat{z}$.
-Since the partial derivative is non-negative, we will obtain a larger utility if we increase the value of $c^{*}(\hat{z})$.
-Hence, when the allocated RB is $\alpha\hat{z}$, the optimal computation workload is larger than $c^{*}(\hat{z})$, i.e.
-$c^{*}(\alpha\hat{z})\geq c^{*}(\hat{z})$.
-Since the value of $\alpha$ is arbitrary, we can conclude that $c^{*}(\hat{z})$ is non-decreasing with respect to $\hat{z}$, which completes the proof.
+In this paper, we have studied the joint optimization of user association and resource allocation for ASC in 5G and beyond networks.
+We discussed the feasibility of ASC and formulated the utility maximization problem under limited energy and delay.
+An efficient three-stage algorithm has been proposed by leveraging the structural characteristics of each subproblem and their interconnections.
+Numerical results indicated that the performance of SemCom can significantly outperform the traditional communication paradigm if we choose appropriate control decisions.